### v1

# bibliography.bib

Abstract
We prove that the space $\mathcal{H}_{\infty}$ of framed infinite volume hyperbolic $3$-manifolds is connected but not path connected. Two proofs of connectivity of this space, which is equipped with the geometric topology, are given, each utilizing the density theorem for Kleinian groups. In particular, we decompose $\mathcal{H}_{\infty}$ into a union of leaves, each leaf corresponding to the set of framings on a fixed manifold, and construct a leaf which is dense in $\mathcal{H}_{\infty}$. Examples of paths in $\mathcal{H}_{\infty}$ are discussed, and machinery for analyzing general paths is developed, allowing for a description of paths of convex cocompact framed manifolds. The discussion of paths culminates in describing an infinite family of non-tame hyperbolic $3$-manifolds whose corresponding leaves are path components of $\mathcal{H}_{\infty}$, which establishes that $\mathcal{H}_{\infty}$ is not path connected.
1 Introduction
We will examine questions of global connectivity in the space $\mathcal{H}$ of isometry classes of complete oriented connected framed hyperbolic $3$-manifolds $(M,f)$. This space is equipped with the geometric topology. Informally, two framed manifolds are close in the geometric topology if they are almost isometric on large compact neighborhoods of their framed basepoints. The space $\mathcal{H}$ is homeomorphic to the space $\mathcal{D}$ of Kleinian groups equipped with the Chabauty topology, which we will also refer to as the geometric topology; here, a Kleinian group is a torsion free discrete subgroup of $\text{PSL}_{2}\mathbb{C}$ (see Section 2.2).
A sequence of Kleinian groups $\{\Gamma_{n}\}\subseteq\mathcal{D}$ converges to $\Gamma$ in the geometric topology on $\mathcal{D}$ if and only if the following two conditions hold:
  1. 1.
    If $\psi\in\text{PSL}_{2}\mathbb{C}$ is an accumulation point of a sequence $\psi_{n}\in\Gamma_{n}$, then $\psi\in\Gamma$.
  2. 2.
    For all $\psi\in\Gamma$, there exists a sequence $\psi_{n}\in\Gamma_{n}$ so that $\psi_{n}\rightarrow\psi$ in $\text{PSL}_{2}\mathbb{C}$.
The resolution of the density conjecture (now theorem) by Namazi and Souto [NAM] and separately Ohshika [OHS] provides a starting place to study the global topology of the space $\mathcal{D}$ of Kleinian groups. A strong version of the density theorem [NAM, Corollary 12.3] says that every finitely generated Kleinian group is the geometric limit of a sequence of geometrically finite Kleinian groups. Any Kleinian group is the geometric limit of a sequence of finitely generated subgroups of itself (see Lemma 2.5), so we have the following corollary of the density theorem.
Density (Namazi-Souto [NAM], Ohshika [OHS]).
Every Kleinian group is the geometric limit of a sequence of geometrically finite Kleinian groups.
We will use this result to identify all connected components of $\mathcal{H}$. For a hyperbolic $3$-manifold $M$ with positively oriented orthonormal frame bundle $\mathcal{F}M$, we define the map $L_{M}:\mathcal{F}M\rightarrow\mathcal{H}$ by $L_{M}(f)=(M,f)$ and set
$\ell(M)=L_{M}(\mathcal{F}M)=\{(M,f)\;|\;f\in\mathcal{F}M\}$
to be the leaf of $\mathcal{H}$ corresponding to $M$.
Theorem A.
The connected components of $\mathcal{H}$ are:
  1. 1.
    $\ell(M)$ for each hyperbolic $3$-manifold $M$ with $\text{vol}(M)<\infty$
  2. 2.
    =$\mathcal{H}_{\infty}=\{(M,f)\in\mathcal{H}\;|\;\text{vol}(M)=\infty\}$.
The fact that each $\ell(M)$ is a connected component of $\mathcal{H}$ for $M$ with $\text{vol}(M)<\infty$ follows from Mostow-Prasad rigidity and classical results on sequences in the geometric topology (see eg. [BEN, Section E]). To determine that the subset $\mathcal{H}_{\infty}$ of infinite volume framed hyperbolic $3$-manifolds is connected, we first show that all infinite volume geometrically finite framed hyperbolic $3$-manifolds lie in the same path component of $\mathcal{H}_{\infty}$ as $(\mathbb{H}^{3},\mathcal{O})$, where $\mathcal{O}\in\mathcal{F}\mathbb{H}^{3}$ (see Lemma 3.3). The density theorem implies that this path component is dense in $\mathcal{H}_{\infty}$, so the closure of this path component must be $\mathcal{H}_{\infty}$, which is therefore connected.
The strategy to show that $\mathcal{H}_{\infty}$ is connected was to construct a path connected subset of $\mathcal{H}$ which is dense in $\mathcal{H}_{\infty}$. Each leaf $\ell(M)$ is path connected, so the following theorem produces another dense path connected subset of $\mathcal{H}_{\infty}$, providing a second proof that $\mathcal{H}_{\infty}$ is connected.
Theorem B.
There exists a hyperbolic $3$-manifold $M$ such that the leaf $\ell(M)$ is dense in $\mathcal{H}_{\infty}$.
This theorem is established by modifying a circle packing construction by Fuchs, Purcell, and Stewart [PUR] (see also Brooks [BRO]) to show that every infinite volume framed hyperbolic $3$-manifold is the geometric limit of a sequence of framed hyperbolic $3$-manifolds whose convex core boundary is a disjoint union of totally geodesic thrice punctured spheres (see Proposition 3.8). These manifolds are then glued together along totally geodesic thrice punctured spheres in their convex core boundaries to construct the manifold $M$ in Theorem B.
Paths in the geometric topology are more controlled than sequences. Understanding this control is one of our major goals. The following proposition (see Section 4.2) provides a way of tracking elements of $\text{PSL}_{2}\mathbb{C}$ through a path in $\mathcal{D}$, and is a key tools for studying paths in $\mathcal{D}$.
Proposition.
Let $\Gamma:I\rightarrow\mathcal{D}$ be a path, and set $\Gamma(t)=\Gamma_{t}$ for all $t$ in the interval $I$. Then, for all $s\in I$, there exists a unique family of maps $J_{s,t}:\Gamma_{s}\rightarrow\overline{\text{PSL}_{2}\mathbb{C}}=\text{PSL}_{2}\mathbb{C}\cup\infty$ for $t\in I$ such that for all $\psi\in\Gamma_{s}$,
  1. 1.
    $J_{s,t}(\psi)\in\Gamma_{t}\cup\infty$
  2. 2.
    $J_{s,s}(\psi)=\psi$
  3. 3.
    the map $t\mapsto J_{s,t}(\psi)\in\overline{\text{PSL}_{2}\mathbb{C}}$ is continuous for $t\in[0,1]$
  4. 4.
    if $s<t^{\prime}<t$ or $t<t^{\prime}<s$, then $J_{s,t^{\prime}}(\psi)=\infty$ implies $J_{s,t}(\psi)=\infty$.
Moreover, for any $H\leq\Gamma_{s}$, $J_{s,t}|_{H}$ is an injective homomorphism into $\text{PSL}_{2}\mathbb{C}$ if $\infty\not\in J_{s,t}(H)$.
The continuity in property (3) of this proposition tells us that any element $\psi\in\Gamma_{s}$ can be tracked through the groups $\Gamma_{t}$ by the path $J_{s,t}(\psi)$ for $t$ close to $s$, and the only way to “lose track” of an element is for this path to diverge to infinity. One can use the Klein-Maskit combination theorems (see eg. [MAT]) to construct examples of paths in $\mathcal{D}$ exhibiting this behavior where paths $t\mapsto J_{s,t}(\psi)$ diverge to infinity.
Example C.2: For Kleinian groups $G$ and $H$ with $\Omega(G),\Omega(H)\neq\varnothing$, there exists a path $\varphi:[0,\infty)\rightarrow\text{PSL}_{2}\mathbb{C}$ so that if $H_{t}=\varphi(t)H\varphi(t)^{-1}$, then the following hold:
  1. 1.
    each $\langle G,H_{t}\rangle$ is discrete and splits as a free product $\langle G,H_{t}\rangle=G*H_{t}$,
  2. 2.
    $t\mapsto G*H_{t}$ is a path in $\mathcal{D}$,
  3. 3.
    $G*H_{t}\rightarrow G$ in $\mathcal{D}$ as $t\rightarrow\infty$.
Example C.2 constructs a path $\Gamma:[0,\infty]\rightarrow\mathcal{D}$ where $\Gamma(t)=\Gamma_{t}=\langle G,H_{t}\rangle$ for $t<\infty$ and $\Gamma(\infty)=G$. Then, for each $\psi\in H_{0}$, the path $t\mapsto J_{0,t}(\psi)=\varphi(t)\psi\varphi(t)^{-1}$ diverges to $\infty$ as $t\rightarrow\infty$. Additionally, $J_{\infty,0}(G)=G$ is a free factor of $\Gamma_{0}$. The next theorem will tell us that the splitting of groups in Example C.2 is in fact reflective of a general phenomenon for paths of convex cocompact Kleinian groups: only free factors can diverge to infinity in a path of convex cocompact Kleinian groups.
Theorem C.
Suppose $\Gamma:[0,1]\rightarrow\mathcal{D}$ is a path such that $\Gamma(t)=\Gamma_{t}$ is convex cocompact for all $t$. Then, $H^{J}=J_{1,0}(\Gamma_{1})\cap\text{PSL}_{2}\mathbb{C}$ is a free factor of $\Gamma_{0}$. In particular, there exist finitely many subgroups $H_{1},...,H_{n}\leq\Gamma_{0}$ and times $0<t_{1}<...<t_{n}\leq 1$ so that
  1. 1.
    $\Gamma_{0}=H^{J}*H_{1}*...*H_{n}$ and
  2. 2.
    $J_{0,t}(\psi)\rightarrow\infty$ as $t\rightarrow t_{i}$, for all $\psi\in H_{i}$ and $1\leq i\leq n$.
Note that the conclusion of Theorem C greatly differs from the analogous situation of a convergent sequence of convex cocompact Kleinian groups. In fact, any Kleinian group $\Gamma$ (in particular, any convex cocompact Kleinian group) may be approximated by a sequence of convex cocompact Kleinian groups which do not split as a non-trivial free product (see Section 4.3). Additionally, the conclusion of Theorem C does not hold if the hypothesis “convex cocompact” is replaced with “geometrically finite” (see Example B.2).
We conclude with a discussion of path connectivity in the space $\mathcal{H}_{\infty}$ of infinite volume framed hyperbolic $3$-manifolds. We define an infinite family of non-tame hyperbolic $3$-manifolds, which we call symmetric infinite type $(G,N)$-glued hyperbolic $3$-manifolds. Imprecisely, manifolds in this family are constructed by gluing together homeomorphic copies of a compact, oriented, connected, irreducible, atoriodal, acylindrical, boundary incompressible $3$-manifold $N$ according to vertex adjacencies in a highly symmetric infinite graph $G$ (see Section 4.4 for a precise definition). These manifolds are hyperbolizable by a theorem of Souto and Stover [SOUSTO]. In fact, the symmetry of $G$ implies that each symmetric infinite type $(G,N)$-glued hyperbolic $3$-manifold has a unique hyperbolic structure, by a forthcoming result of Cremaschi and Yarmola [CREMYAR] (see Proposition 4.16). Our main result is the following.
Theorem D.
For any symmetric infinite type $(G,N)$-glued hyperbolic $3$-manifold $M$, the leaf $\ell(M)$ is a path component of $\mathcal{H}_{\infty}$.
Roughly, the proof proceeds by supposing that $\Gamma:[0,1]\rightarrow\mathcal{H}$ is a path so that $M=\mathbb{H}^{3}/\Gamma(0)$ is a symmetric infinite type $(G,N)$-glued hyperbolic $3$-manifold. Then, one considers a sequence of subgroups $H_{n}\leq H_{n+1}\leq\Gamma(0)$ exhausting $\Gamma(0)$ corresponding to an exhaustion $M=\cup X_{n}$, where each $X_{n}$ is a connected union of copies of $N$. If the image of any map $J_{0,t}|_{H_{n}}$ contains $\infty$, then one obtains a contradiction to Thurston’s “$AH(\text{acylindrical})$ is compact” theorem [THUI], using that each $X_{n}$ is acylindrical. One then concludes that $J_{0,t}$ must be an injective homomorphism for all $t$, so the rigidity of $M$ implies that each $\Gamma(t)$ contains a subgroup which is conjugate to $\Gamma(0)$. Finally, a result on discrete extensions of Kleinian groups (Proposition 4.18) implies that each $\Gamma(t)$ itself must in fact be conjugate to $\Gamma(0)$.
Theorem D has the following immediate corollary.
Corollary E.
$\mathcal{H}_{\infty}$ is not path connected.
Our analysis here combines elements of the classical theory of deformations of infinite volume hyperbolic $3$-manifolds with the study of Chabauty spaces of subgroups of Lie groups. Questions similar to ours have been been considered in each of these settings. In particular, Warakkagun [WAR] has examined the two-dimensional version of our setting, proving that the Chabauty space of torsion free discrete subgroups of $\text{PSL}_{2}\mathbb{R}$ is path connected, in direct contrast to Corollary E. Additionally, Baik and Clavier [BAIK], have studied geometric limits of cyclic subgroups of $\text{PSL}_{2}\mathbb{R}$, later generalizing to geometric limits of abelian subgroup of $\text{PSL}_{2}\mathbb{C}$ [BAIK2]. Biringer, Lazarovich, and Leitner [IANCHAB] have further analyzed the space of closed subgroups of $\text{PSL}_{2}\mathbb{R}$, focusing on global topology. We also note that a result of Fraczyk and Gelander [FRAGEL] implies that the space of discrete subgroups of $\text{SL}_{n}\mathbb{R}$ for $n\geq 3$ which act with infinite covolume on the symmetric space $\text{SL}_{n}\mathbb{R}/\text{SO}(n)$ is connected, providing a result analogous to Theorem A.
The classical study of deformation spaces of infinite volume hyperbolic $3$-manifolds focuses on the space $AH(\Gamma)$ of conjugacy classes of discrete and faithful representations of a Kleinian group $\Gamma$ into $\text{PSL}_{2}\mathbb{C}$ (see Section 2.3). Sullivan’s rigidity theorem [SUL] identifies the components of the interior of $AH(\Gamma)$. Anderson and Canary [ACBumping] show that the intersection of closures of components of the interior of $AH(\Gamma)$ may be non-empty in a phenomenon called self-bumping, which they study more thoroughly with McCullough in [ACM]. Bromberg [KEN], and later Brock, Bromberg, Canary, and Minsky [BBCM], study local connectivity in $AH(\Gamma)$.
We now describe the organization of the rest of the paper. In Section 2 we give background information on Kleinian groups, hyperbolic $3$-manifolds, the geometric topology, and spaces of representations of Kleinian groups. Section 3.1 focuses on determining the connected components of $\mathcal{H}$, in particular proving Theorem A. Section 3.2 constructs a manifold $M$ so that the leaf $\ell(M)$ is dense in $\mathcal{H}_{\infty}$, proving Theorem B. In Section 4, we turn our attention to paths. Examples of paths are produced in Section 4.1 and the basic machinery for discussing paths in $\mathcal{D}$ is developed in Section 4.2, in particular constructing the maps $J_{s,t}$. Theorem C on paths of convex cocompact Kleinian groups is proved in Section 4.3. Finally, we conclude with Section 4.4, in which we define $(G,N)$-glued hyperbolic $3$-manifolds and prove Theorem D, establishing Corollary E that $\mathcal{H}_{\infty}$ is not path connected.
Acknowledgements: The author would like to start by thanking Ian Biringer for numerous conversations and suggestions, and for his unending kindness. The author thanks Tommaso Cremaschi and Andrew Yarmola for allowing him to include a result from a forthcoming paper [CREMYAR], appearing here as Proposition 4.16. The author also thanks Cremaschi for helpful conversations.
2 Background
Throughout the paper, all manifolds considered are connected and oriented. Any hyperbolic manifold will be complete unless otherwise noted.
2.1 Kleinian groups and hyperbolic 3-manifolds
Identifying the Riemann sphere with the ideal boundary $S^{2}_{\infty}$ of $\mathbb{H}^{3}$, Poincaré extensions of Möbius transformations allow us to identify $\text{PSL}_{2}\mathbb{C}$ and $\text{Isom}^{+}(\mathbb{H}^{3})$, the group of orientation preserving isometries of hyperbolic $3$-space. A Kleinian group is a discrete and torsion free subgroup of $\text{PSL}_{2}\mathbb{C}$; note that in many references a Kleinian group is not required to be torsion free, but we include this in our definition. We let $\mathcal{D}$ denote the set of Kleinian groups:
$\mathcal{D}=\{\Gamma\leq\text{PSL}_{2}\mathbb{C}\;|\;\Gamma\text{ is discrete and torsion free}\}.$
A hyperbolic $3$-manifold $M$ is the quotient $\mathbb{H}^{3}/\Gamma$ for some Kleinian group $\Gamma$. From $\Gamma$, $M$ inherits a $(\text{PSL}_{2}\mathbb{C},\mathbb{H}^{3})$-structure, in the language of $(G,X)$-structures (see [THUR], [BEN]). For any $p\in M$, we obtain a holonomy representation of $\pi_{1}(M)$ with respect to $p$ and $\Gamma$, an isomorphism $\Psi:\pi_{1}(M,p)\rightarrow\Gamma$. The holonomy representation $\Psi$ is unique up to conjugation by $\Gamma$, and is determined by choosing a $\Gamma$-lift of $p$ to $\mathbb{H}^{3}$. Hyperbolic $3$-manifolds will be identified if there is an orientation preserving isometry between them.
For a Kleinian group $\Gamma$, the limit set $\Lambda(\Gamma)$ of $\Gamma$ is constructed by selecting a point $x\in\mathbb{H}^{3}$ and letting $\Lambda(\Gamma)=\overline{\Gamma\cdot x}\cap S^{2}_{\infty}$. This definition is independent of the choice of $x$. Additionally, define the domain of discontinuity of $\Gamma$ as $\Omega(\Gamma)=S^{2}_{\infty}\backslash\Lambda(\Gamma)$, and recall that $\Omega(\Gamma)$ is the largest subset of $S^{2}_{\infty}$ on which $\Gamma$ acts properly discontinuously.
Let $CH(\Gamma)\subseteq\mathbb{H}^{3}$ be the (hyperbolic) convex hull of $\Lambda(\Gamma)$. Then, for the hyperbolic $3$-manifold $M=\mathbb{H}^{3}/\Gamma$, we define the convex core of $M$ to be $CC(M)=CH(\Gamma)/\Gamma\subseteq M$. $M$ and $\Gamma$ are each called geometrically finite if $\Gamma$ is finitely generated and $CC(M)$ has finite volume; $M$ and $\Gamma$ are called convex cocompact if $CC(M)$ is compact. Finally, $M$ and $\Gamma$ are said to be elementary if $|\Lambda(\Gamma)|\leq 2$, which is equivalent to $\Gamma$ being abelian. See [NOTE] and [MAT] for more background on Kleinian groups.
A subset $U$ of either $S^{2}_{\infty}$ or $\mathbb{H}^{3}$ is said to be precisely invariant for a subgroup $H\leq\Gamma$ if $\psi(U)=U$ for all $\psi\in H$ and $\psi(U)\cap U=\varnothing$ for all $\psi\in\Gamma\backslash H$. For $p\in M=\mathbb{H}^{3}/\Gamma$, the injectivity radius of $M$ at $p$ is
$\text{inj}_{M}(p)=\sup\{R>0:B_{\mathbb{H}^{3}}(\tilde{p},R)\text{ is preciselyinvariant for }\{\mathds{1}\}\text{ in }\Gamma\}$
where $\tilde{p}\in\mathbb{H}^{3}$ is a $\Gamma$-lift of $p$ and $\mathds{1}\in\text{PSL}_{2}\mathbb{C}$ is the identity. In other words, $\text{inj}_{M}(p)$ is the supremal radius of hyperbolic balls centered at $p\in M$ which isometrically embed into $M$. For any $\varepsilon>0$, we set $M_{\leq\varepsilon}=\{p\in M\;|\;\text{inj}_{M}(p)\leq 2\varepsilon\}$ to be the $\varepsilon$-thin part of $M$.
A frame $f$ for a hyperbolic $3$-manifold $M$ at $p\in M$ is an ordered positively oriented orthonormal basis for the tangent space $T_{p}M$. For a fixed $p\in M$, let $F_{p}M$ be the space of frames for $M$ at $p$. When we would like to make explicit the basepoint underlying a frame in $F_{p}M$, we write the frame as $f_{p}$. Let $\mathcal{F}M$ be the bundle of positively oriented orthonormal frames over $M$. As a set, for any $U\subseteq M$, in particular when $U=M$,
$\mathcal{F}U=\bigcup_{p\in U}F_{p}M.$
A pair $(M,f)$ where $f\in\mathcal{F}M$ is a framed hyperbolic $3$-manifold. Let
=$\mathcal{H}=\{(M,f)\;|\;M\text{ hyperbolic }3\text{-manifold},f\in\mathcal{F}M\}/\sim$
where $(M,f)\sim(N,h)$ if there is an isometry $\varphi:M\rightarrow N$ such that the induced map $\varphi_{*}:\mathcal{F}M\rightarrow\mathcal{F}N$ satisfies $\varphi_{*}(f)=h$. We will refer to equivalence classes $[(M,f)]\in\mathcal{H}$ by their representatives, for example $(M,f)$. Background on $\mathcal{H}$ can be found in [BEN].
Once and for all, fix $O\in\mathbb{H}^{3}$ and $\mathcal{O}_{O}\in F_{O}\mathbb{H}^{3}$. We can now define a map $\Phi:\mathcal{D}\rightarrow\mathcal{H}$ by
$\Phi(\Gamma)=(\mathbb{H}^{3}/\Gamma,\pi_{\Gamma}(\mathcal{O}_{O}))$
where $\pi_{\Gamma}:\mathcal{F}\mathbb{H}^{3}\rightarrow\mathcal{F}(\mathbb{H}^{3}/\Gamma)$ is the natural projection. As discussed in [BEN, E.1.9], $\Phi$ is in fact a bijection, which is a primary reason for working with framed hyperbolic $3$-manifolds, rather than, say, pointed hyperbolic $3$-manifolds.
2.2 The geometric topology
We will describe a topology on $\mathcal{H}$. First, we make the following definition:
Definition 2.1.
Let $(M,f_{p})=(\mathbb{H}^{3}/\Gamma_{1},\pi_{\Gamma_{1}}(\mathcal{O}_{O}))$ and $(N,h_{q})=(\mathbb{H}^{3}/\Gamma_{2},\pi_{\Gamma_{2}}(\mathcal{O}_{O}))$ be framed hyperbolic $3$-manifolds. We say $(N,h_{q})$ is $(\varepsilon,R)$-close to $(M,f_{p})$ if there is a $(1+\varepsilon)$-bilipshitz embedding $\tilde{g}:B_{\mathbb{H}^{3}}(O,R)\rightarrow\mathbb{H}^{3}$ such that
  1. 1.
    $\tilde{g}(O)=O$
  2. 2.
    $D_{C^{0}}(\tilde{g},\mathds{1}|_{B_{\mathbb{H}^{3}}(O,R)})<\varepsilon$
  3. 3.
    $\tilde{g}$ descends to an embedding $g:B_{M}(p,R)\rightarrow N$.
Here, for $U\subseteq\mathbb{H}^{3}$ and $g_{1},g_{2}:U\rightarrow\mathbb{H}^{3}$,
$D_{C^{0}}(g_{1},g_{2})=\sup_{z\in U}d_{\mathbb{H}^{3}}(g_{1}(z),g_{2}(z)).$
Definition 2.2.
The geometric topology on $\mathcal{H}$ is the topology generated by taking the collection of sets of the form
$\{(N,q)\in\mathcal{H}\;|\;(N,q)\text{ is }(\varepsilon,R)\text{-close to }(M,p)\},$
where $(M,p)\in\mathcal{H}$ and $\varepsilon,R>0$ as a subbasis.
Similar, definitions of $(\varepsilon,R)$-close are given in [BEN], [NOTE], and [PUR]. This topology is metrizable (see [BEN, E.1.4] and [IANMET]). The following lemma provides a type of transitivity for being $(\varepsilon,R)$-close, and follows immediately from Definition 2.1.
Lemma 2.3.
Suppose $(N,h_{q})$ is $(\varepsilon,R)$-close to $(M,f)$, and that for some hyperbolic $3$-manifold $N^{\prime}$, there exists an isometric embedding $\iota:B_{N}(q,R)\rightarrow N^{\prime}$. Then, $(N^{\prime},\iota_{*}(h))$ is $(\varepsilon,R)$-close to $(M,f)$, where $\iota_{*}:\mathcal{F}B_{N}(q,R)\rightarrow\mathcal{F}N^{\prime}$ is the induced map on frame bundles.
If $\mathcal{D}$ is equipped with the topology inherited from the Chabauty topology (see [BEN, Section E]) on closed subsets of $\text{PSL}_{2}\mathbb{C}$, the map $\Phi$ is a homeomorphism. For clarity, we will therefore also refer to the Chabauty topology on $\mathcal{D}$ as the geometric topology. The following fact (see [BEN, Proposition E.1.2]) describes sequential convergence in the geometric topology on $\mathcal{D}$.
Fact 2.4.
A sequence $\{\Gamma_{n}\}\subseteq\mathcal{D}$ converges to $\Gamma$ in the geometric topology on $\mathcal{D}$ if and only if the following two conditions hold:
  1. 1.
    If $\psi\in\text{PSL}_{2}\mathbb{C}$ is an accumulation point of a sequence $\psi_{n}\in\Gamma_{n}$, then $\psi\in\Gamma$.
  2. 2.
    For all $\psi\in\Gamma$, there exists a sequence $\psi_{n}\in\Gamma_{n}$ so that $\psi_{n}\rightarrow\psi$ in $\text{PSL}_{2}\mathbb{C}$.
The map $\Phi:\mathcal{D}\rightarrow\mathcal{H}$ is a homeomorphism between the geometric topologies on these sets.
Note that for a convergent sequence $\Gamma_{n}\rightarrow\Gamma$, the groups $\Gamma_{n}$ are not required to be isomorphic to each other or to $\Gamma$. Often, the geometric topology is discussed in the context of sequences of representations of a fixed group, but this is not the case here.
It will be useful to have the following lemma recorded.
Lemma 2.5.
Any Kleinian group is the geometric limit of a sequence of finitely generated subgroups of itself.
Proof.
For a Kleinian $\Gamma$, enumerate $\Gamma=\{\psi_{1},\psi_{2},...\}$ and let $\Gamma_{n}=\langle\psi_{1},...,\psi_{n}\rangle$. We will show that the sequence $\{\Gamma_{n}\}$ converges geometrically to $\Gamma$.
We just need to check the two conditions in Fact 2.4. Observe that the first condition holds since $\Gamma$ is a closed subset of $\text{PSL}_{2}\mathbb{C}$ and $\Gamma_{n}\subseteq\Gamma$ for all $n$. The second condition is immediate, since any $\psi\in\Gamma$ is contained in $\Gamma_{n}$ for $n$ sufficiently large. ∎
2.3 Discrete and faithful representations
Let $\Gamma$ be a nonelementary Kleinian group. An injective homomorphism $\rho:\Gamma\rightarrow\text{PSL}_{2}\mathbb{C}$ whose image is a Kleinian group is called a discrete and faithful representation of $\Gamma$ into $\text{PSL}_{2}\mathbb{C}$, and we let $D(\Gamma)\subseteq\text{Hom}(\Gamma,\text{PSL}_{2}\mathbb{C})$ denote the set such representations. Equip $D(\Gamma)$ with the algebraic topology in which a sequence of representations $\{\rho_{n}\}\subseteq D(\Gamma)$ algebraically converges to $\rho\in D(\Gamma)$ if
$\lim_{n}\rho_{n}(\psi)=\rho(\psi)$
for all $\psi\in\Gamma$, where convergence of the limit is in $\text{PSL}_{2}\mathbb{C}$. An algebraically convergent sequence $\rho_{n}\rightarrow\rho$ in $D(\Gamma)$ is said to converge strongly if the images of the representations converge geometrically: that is, $\rho_{n}(\Gamma)\rightarrow\rho(\Gamma)$ geometrically in $\mathcal{D}$. We let $S(\Gamma)$ denote the set $D(\Gamma)$ now equipped with the topology of strong convergence.
Let $AH(\Gamma)=D(\Gamma)/\text{conj}$ denote the quotient of $D(\Gamma)$ by the conjugation action of $\text{PSL}_{2}\mathbb{C}$, and equip $AH(\Gamma)$ with the topology inherited from the algebraic topology on $D(\Gamma)$. This topology on $AH(\Gamma)$ is also referred to as the algebraic topology. Note that a sequence $\{[\rho_{n}]\}\subseteq AH(\Gamma)$ algebraically converges to $[\rho]$ if and only if there exists a sequence $\{\psi_{n}\}\subseteq\text{PSL}_{2}\mathbb{C}$ such that $\psi_{n}\rho_{n}\psi_{n}^{-1}$ converges to $\rho$ in $D(\Gamma)$. See [MAT] for more details on the algebraic topology.
A representation $\rho\in D(\Gamma)$ is type preserving if $\rho(\psi)$ is parabolic exactly when $\psi\in\Gamma$ is parabolic, and we let $D_{t}(\Gamma)\subseteq D(\Gamma)$ be the subspace of type preserving representations. If $\Gamma$ is geometrically finite, we let $GF(\Gamma)$ denote the component of the interior of $D_{t}(\Gamma)$ containing the identity mapping. It follows from the Marden [MAR] and Sullivan stability theorems [SUL] that $GF(\Gamma)$ consists of all representations of $\Gamma$ whose image is a geometrically finite Kleinian group which is quasiconformally conjugate to $\Gamma$: that is, there exists a quasiconformal homeomorphism $F:S^{2}_{\infty}\rightarrow S^{2}_{\infty}$ such that for all $\psi\in\Gamma$, $\rho(\psi)=F\circ\psi\circ F^{-1}$. If $\Gamma$ is convex cocompact, the image of every representation in $GF(\Gamma)$ is convex cocompact. We let $\overline{GF}(\Gamma)=GF(\Gamma)/\text{conj}$ denote the projection of $GF(\Gamma)$ to $AH(\Gamma)$.
We will record a general fact about representations of nonelementary Kleinian groups, which follows from the fact that $\text{PSL}_{2}\mathbb{C}$ acts uniquely triply transitively on $S^{2}_{\infty}$.
Fact 2.6.
Suppose $\rho_{1},\rho_{2}\in D(\Gamma)$ represent the same conjugacy class in $AH(\Gamma)$ for some nonelementary Kleinian group $\Gamma$. Then, there is a unique $\varphi\in\text{PSL}_{2}\mathbb{C}$ such that $\rho_{2}=\varphi\rho_{1}\varphi^{-1}$. Additionally, if $\rho_{1},\rho_{2}$ are perturbed in $D(\Gamma)$ while remaining conjugate, then $\varphi$ varies continuously.
The continuity claim at the end of the fact follows from writing $\varphi$ as a continuous function of the fixed points of some elements of each $\rho_{i}(\Gamma)$, the fixed points being continuous functions of the isometries themselves. From this fact, we can prove the following lemma.
Lemma 2.7.
For a Kleinian group $\Gamma$, let $[\rho_{n}]\rightarrow[\rho]$ be a convergent sequence in $AH(\Gamma)$. Assume that there exists a nonelementary subgroup $H\leq\Gamma$ and a representative $\rho_{n}\in D(\Gamma)$ of each conjugacy class $[\rho_{n}]$ so that for some $\rho_{H}\in D(H)$, we have $\rho_{n}|_{H}\rightarrow\rho_{H}$ in $D(H)$. Then, there exists a representative $\rho$ of $[\rho]$ such that $\rho_{n}\rightarrow\rho$ in $D(\Gamma)$. In particular, $\rho|_{H}=\rho_{H}$.
Proof.
Let $\{\psi_{n}\}\subseteq\text{PSL}_{2}\mathbb{C}$ be a sequence so that $\psi_{n}\rho_{n}\psi_{n}^{-1}$ converges in $D(\Gamma)$ to some representation $\rho^{\prime}$. Since the sequence $\{\rho_{n}|_{H}\}$ converges, the sequence $\{\psi_{n}\}$ must be bounded in $\text{PSL}_{2}\mathbb{C}$, and therefore the sequence $\{\rho_{n}\}$ must be bounded in $D(\Gamma)$. A theorem of Jørgensen [JOR] tells us $D(\Gamma)$ is closed in $\text{Hom}(\Gamma,\text{PSL}_{2}\mathbb{C})$, so it follows that for any subsequence of $\{\rho_{n}\}$, there exists a further subsequence that converges to some $\rho\in D(\Gamma)$. Observe that this representation $\rho$ is independent of the initial subsequence chosen: $\rho$ must be conjugate in $\text{PSL}_{2}\mathbb{C}$ to $\rho^{\prime}$, and for each initial subsequence it follows that $\rho|_{H}=\rho_{H}$, so the observation follows from the uniqueness statement in Fact 2.6. Since every subsequence of $\{\rho_{n}\}$ has a subsequence which converges to $\rho$, we have $\rho_{n}\rightarrow\rho$. ∎
3 Connectivity
This section will focus on connectivity in $\mathcal{H}$. Section 3.1 will focus on establishing that the subspace =$\mathcal{H}_{\infty}=\{(M,f)\;|\;\text{vol}(M)=\infty\}$ is connected, and in Section 3.2 we will construct a non-tame hyperbolic $3$-manifold $M$ such that the leaf $\ell(M)$ (see Section 3.1) is dense in $\mathcal{H}_{\infty}$.
3.1 Connected Components
The goal of this section will be to determine the connected components of $\mathcal{H}$ and in particular show that the subspace $\mathcal{H}_{\infty}$ is connected. Briefly, the idea will be to use the density theorem to show that there is a path connected subset which is dense in $\mathcal{H}_{\infty}$.
For a hyperbolic $3$-manifold $M$, we have a map $L_{M}:\mathcal{F}M\rightarrow\mathcal{H}$ such that $L_{M}(f)=(M,f)$. The set
$\ell(M)=L_{M}(\mathcal{F}M)=\{(M,f)\in\mathcal{H}\;|\;f\in\mathcal{F}M\}$
is called the leaf of $\mathcal{H}$ corresponding to $M$. Biringer and Abért discuss a similar leaf decomposition of the space of pointed Riemannian $d$-manifolds in [IAN].
Following from the equivalence relation defining $\mathcal{H}$, $L_{M}(f)=L_{M}(h)$ if and only if there is an orientation preserving isometry of $M$ such that the induced map on $\mathcal{F}M$ takes $f$ to $h$. Thus, there is a bijection $\ell(M)\leftrightarrow\mathcal{F}M/\text{Isom}^{+}(M)$ descending from $L_{M}$.
Proposition 3.1.
$L_{M}$ is continuous for any hyperbolic $3$-manifold $M$.
Proof.
It will suffice to show that the map $\Phi^{-1}\circ L_{M}:\mathcal{F}M\rightarrow\mathcal{D}$ is continuous, where $\Phi:\mathcal{D}\rightarrow\mathcal{H}$ is the homeomorphism from Section 2.1. Fix a convergent sequence $f_{n}\rightarrow f$ in $\mathcal{F}M$ and set $\Gamma=\Phi^{-1}((M,f))$. Let $\{\tilde{f}_{n}\}\subseteq\mathcal{F}\mathbb{H}^{3}$ be a $\Gamma$-lift of the sequence $\{f_{n}\}$ so that $\tilde{f}_{n}\rightarrow\mathcal{O}_{O}$ as $n\rightarrow\infty$. Recalling that $\text{PSL}_{2}\mathbb{C}$ acts simply transitively on $\mathcal{F}\mathbb{H}^{3}$, set $\psi_{n}\in\text{PSL}_{2}\mathbb{C}$ so that $\psi_{n}(\tilde{f}_{n})=\mathcal{O}_{O}$. Then, we have $\psi_{n}\rightarrow\mathds{1}$ as $n\rightarrow\infty$, so
$(\Phi^{-1}\circ L_{M})(f_{n})=\psi_{n}\Gamma\psi_{n}^{-1}\rightarrow\Gamma=(\Phi^{-1}\circ L_{M})(f)$
as $n\rightarrow\infty$. ∎
Each frame bundle $\mathcal{F}M$ is path connected, so we have the following immediate corollary.
Corollary 3.2.
For any hyperbolic $3$-manifold $M$, $\ell(M)$ is path connected in $\mathcal{H}$.
Lemma 3.3.
All framed hyperbolic $3$-manifolds $(\mathbb{H}^{3}/\Gamma,f)\in\mathcal{H}$ such that $\Omega(\Gamma)\neq\varnothing$ are contained in the same path component of $\mathcal{H}$.
Proof.
Let $M=\mathbb{H}^{3}/\Gamma$, where $\Omega(\Gamma)\neq\varnothing$. We will show that $(M,f)$ is in the same path component of $\mathcal{H}$ as $(\mathbb{H}^{3},\mathcal{O}_{O})$. Pick $x\in\Omega(\Gamma)$. Since $\Omega(\Gamma)$ is open and $\Gamma$ acts properly discontinuously on $\Omega(\Gamma)$, we may pick an open circular disk $D\subseteq\Omega(\Gamma)$ containing $x$ such that $D$, and hence the convex hull $CH(D)\subseteq\mathbb{H}^{3}$, is precisely invariant for $\{\mathds{1}\}$ in $\Gamma$.
Let $\gamma:[0,\infty)\rightarrow\mathbb{H}^{3}$ be a geodesic ray such that $\gamma(0)=O$ and $\gamma(t)$ limits to $x$. Since $CH(D)$ isometrically embeds into $M$, we see that for any $R>0$, $B_{\mathbb{H}^{3}}(\gamma(t),R)$ isometrically embeds into $M$ for $t$ sufficiently large. Thus, letting $f^{t}\in\mathcal{F}M$ be the $\Gamma$-projection of the time $t$ parallel transport of $\mathcal{O}_{O}$ along $\gamma$, Lemma 2.3 implies that $(M,f^{t})$ geometrically converges to $(\mathbb{H}^{3},\mathcal{O}_{O})$ as $t\rightarrow\infty$. ∎
Lemma 3.3 provides a large path component of $\mathcal{H}_{\infty}$. The condition that $\Omega(\Gamma)\neq\varnothing$ occurs for a number of large families of Kleinian groups. In particular, this condition holds if $\Gamma$ is geometrically finite and has infinite covolume.
Lemma 3.4.
Any $(M,f)\in\mathcal{H}_{\infty}$ is the geometric limit of a sequence of geometrically finite elements of $\mathcal{H}_{\infty}$.
Proof.
Namazi and Souto [NAM, Corollary 12.3] proved a strong version of the density theorem which tells us that all finitely generated elements of
$\mathcal{D}_{\infty}=\{\Gamma\in\mathcal{D}\;|\;\text{vol}(\mathbb{H}^{3}/\Gamma)=\infty\}=\Phi^{-1}(\mathcal{H}_{\infty})$
are geometric limits of geometrically finite elements of $\mathcal{D}_{\infty}$. Now, the result follows from Lemma 2.5. ∎
Theorem A.
The connected components of $\mathcal{H}$ are:
  1. 1.
    $\ell(M)$ for each hyperbolic $3$-manifold $M$ with $\text{vol}(M)<\infty$
  2. 2.
    $\mathcal{H}_{\infty}$.
Proof.
Classical results on sequences in the geometric topology (see eg. [BEN, Theorem E.2.4]) tell us that if $\{(M_{i},f^{i})\}\subseteq\mathcal{H}$ is any sequence converging to a finite volume $(M,f)\in\mathcal{H}$, then for $i$ sufficiently large, either $M_{i}$ is isometric to $M$ (in which case $(M_{i},f^{i})\in\ell(M)$) or $M_{i}$ has finite volume and strictly fewer cusps than $M$. Hence, one may proceed by induction on $n$ to show that for any finite volume hyperbolic $3$-manifold $M$ with $n$ cusps, $\ell(M)$ is a connected component of $\mathcal{H}$.
To show that $\mathcal{H}_{\infty}$ is its own connected component, it will now suffice to show that $\mathcal{H}_{\infty}$ is connected: indeed, we can write $\mathcal{H}$ as the disjoint union
$\mathcal{H}=\mathcal{H}_{\infty}\cup\left(\bigcup_{\text{vol}(M)<\infty}\ell(M)\right),$
where we already know that the leaves in the union are their own connected components. Note that Proposition 3.3 tells us all $(M,f)\in\mathcal{H}_{\infty}$ with $M$ geometrically finite are contained in a single path component, and Lemma 3.4 then tells us that this path component is dense in $\mathcal{H}_{\infty}$. Since the closure of a connected set is connected, we conclude that $\mathcal{H}_{\infty}$ is connected, as desired. ∎
3.2 Dense Leaf
In this subsection, we will construct a hyperbolic $3$-manifold $M$ such that the leaf $\ell(M)$ is a dense subset of $\mathcal{H}_{\infty}$. In other words, any infinite volume framed hyperbolic $3$-manifold may be approximated by appropriate framings on $M$. By Corollary 3.2, $\ell(M)$ is consequently another dense path connected subset of $\mathcal{H}_{\infty}$ (cf. Lemma 3.3), providing a second proof that $\mathcal{H}_{\infty}$ is connected. The construction of $M$ will feature a modification of a construction by Fuchs, Purcell and Stewart in [PUR], which utilizes the circle packing machinery developed by Brooks [BRO].
Definition 3.5.
Let $\Gamma\in\mathcal{D}$ be convex cocompact with infinite covolume. A circle packing $P$ on $\partial_{\infty}(\mathbb{H}^{3}/\Gamma)=\Omega(\Gamma)/\Gamma$ is a $\Gamma$-invariant collection of (projective) circles on $\Omega(\Gamma)$ together with a triangulation $V$ of $\Omega(\Gamma)$ with $P=\{c_{v}\;|\;v\text{ a vertex of }V\}$ satisfying the following:
  1. 1.
    The circles of $P$ bound disks with disjoint interiors
  2. 2.
    Each circle $c_{v}$ is centered at the vertex $v$
  3. 3.
    $c_{v},c_{u}\in P$ are tangent if and only if $\langle v,u\rangle$ is an edge of $V$
  4. 4.
    Arcs of mutually tangent circles $c_{v},c_{u},c_{w}\in P$ form a curvilinear triangle in $\Omega(\Gamma)$ if and only if $\langle u,v,w\rangle$ forms a positively oriented face of $V$
  5. 5.
    Any compact subset of $\Omega(\Gamma)$ intersects finitely many circles in $P$.
A framed convex cocompact hyperbolic $3$-manifold $(M,f)$ admits a circle packing if a circle packing can be constructed on $\partial_{\infty}(M)$ with respect to the Kleinian group $\Phi^{-1}((M,f))$.
Figure 1: This depicts a portion of a circle packing $P$ with its associated triangulation. The shaded regions are curvilinear triangles in $\partial_{\infty}(\mathbb{H}^{3}/\Gamma)$. The dotted circles are the dual circles in $P^{*}$, including $c^{(v,u,w)}$.
To elucidate the fourth condition in Definition 3.5, Figure 1 depicts a portion of a circle packing, in which curvilinear triangles (such as that bounded by the circles $c_{v},c_{u}$, and $c_{w}$) are shown shaded. Note that the face $\langle v,z,w\rangle$ in the associated triangulation is not positively oriented, hence these circles do not bound a curvilinear triangle.
Fix a convex cocompact hyperbolic $3$-manifold $M=\mathbb{H}^{3}/\Gamma$ and a circle packing $P$ of $\partial_{\infty}(M)$ with respect to $\Gamma$. For circles $c_{v},c_{u},c_{w}$ in $P$ forming a curvilinear triangle in $\Omega(\Gamma)$, there is a unique dual circle $c^{(v,u,w)}$ orthogonal to each of $c_{v},c_{u}$, and $c_{w}$ and intersecting them at the points of tangency. The dual circles, including $c^{(v,u,w)}$, are depicted dotted in Figure 1. The collection of all such dual circles corresponding to curvilinear triangles will be denoted $P^{*}$ and called the dual circle packing of $P$. Note that $P^{*}$ may not be an actual circle packing in the sense defined above: for example, the dual graph to a triangulation is not necessarily a triangulation.
For each circle $c$ in $P$ or $P^{*}$, let $H(c)\subseteq\mathbb{H}^{3}$ be the open hyperbolic half space meeting $S^{2}_{\infty}$ at the interior of $c$. We then define the scooped manifold
$M_{P}=\left(\mathbb{H}^{3}\;\backslash\bigcup_{c\in P,P^{*}}H(c)\right){{{/}}}\Gamma.$
The following lemma, proved by Fuchs, Purcell and Stewart in [PUR, Prop 3.3], will tell us that that the boundary of a scooped manifold $M_{P}$ consists of hyperbolic ideal polyhedra. The faces of the polyhedra descend from the boundaries of $H(c)\subseteq\mathbb{H}^{3}$ for $c\in P\cup P^{*}$ and their edges descend from intersections of $H(c)$ and $H(c^{*})$ for $c\in P$ and $c^{*}\in P^{*}$. Faces descending from boundaries of $H(c)$ for $c\in P$ will be colored white and faces coming from boundaries of $H(c^{*})$ for $c^{*}\in P^{*}$ will be colored black.
Lemma 3.6.
Let $(M,f)\in\mathcal{H}_{\infty}$ be convex cocompact such that $\partial_{\infty}M$ admits a circle packing. Then, the scooped manifold $M_{P}$ has the following properties:
  1. 1.
    With the coloring above, no two faces of $\partial M_{P}$ of the same color share an edge.
  2. 2.
    The faces consist of totally geodesic ideal polygons, where the black faces are ideal triangles.
  3. 3.
    The dihedral angle between faces is $\pi/2$.
Using the circle packing machinery developed by Brooks [BRO] combined with the density theorem [NAM], Fuchs, Purcell and Stewart also prove the following result [PUR], which tells us that we can approximate any geometrically finite framed hyperbolic $3$-manifold by convex cocompact framed manifolds whose boundaries admit a circle packing.
Lemma 3.7.
For any geometrically finite $(M,f)\in\mathcal{H}_{\infty}$ and $\varepsilon,R>0$, there exists a convex cocompact $(N,h_{q})\in\mathcal{H}_{\infty}$ which is $(\varepsilon,R)$-close to $(M,f)$ and admits a circle packing $P$ with $B_{N}(q,R)$ contained in the scooped manifold $N_{P}$.
We will now utilize Lemmas 3.6 and 3.7 to give a modification of Lemma 3.7 in which we approximate any framed hyperbolic $3$-manifold by ones whose convex core boundaries are given by a disjoint union of totally geodesic thrice punctured spheres. This construction is a modification that carried out by Fuchs, Purcell, and Stewart in [PUR, Construction 3.10], with different goals.
Proposition 3.8.
For any $(M,f)\in\mathcal{H}_{\infty}$ and $\varepsilon,R>0$, there exists a geometrically finite $(N,h_{q})\in\mathcal{H}_{\infty}$ which is $(\varepsilon,R)$-close to $(M,f)$ such that $\partial CC(N)$ is a union of totally geodesic thrice punctured spheres, with $B_{N}(q,R)\subseteq CC(N)$.
Proof.
Lemma 3.4 tells us that it will suffice to assume $M$ is geometrically finite. As in Lemma 3.7, choose a convex cocompact $(N^{\prime},h_{q^{\prime}}^{\prime})\in\mathcal{H}_{\infty}$ such that $(N^{\prime},h_{q^{\prime}}^{\prime})$ is $(\varepsilon,R)$-close to $(M,f)$ and admits a circle packing $P$ with $B_{N^{\prime}}(q^{\prime},R)$ contained in the scooped manifold $N^{\prime}_{P}$.
Let $\overline{N^{\prime}_{P}}$ be a copy of $N^{\prime}_{P}$, with the reversed orientation. Form $N^{\prime\prime}$ by identifying each white face of $N^{\prime}_{P}$ with its copy in $\overline{N^{\prime}_{P}}$, via the identity map between these faces. Observe that $N^{\prime\prime}$ is a hyperbolic $3$-manifold with boundary given by a disjoint union of totally geodesic thrice punctured spheres: indeed, when mirrored white faces are identified, the edges of black triangles are glued with their mirrored copies, with angles along the glued edges adding up to $\pi$, by Lemma 3.6. Note that $\text{vol}(N_{P}^{\prime})<\infty$, so $\text{vol}(N^{\prime\prime})=2\text{vol}(N_{P}^{\prime})<\infty$.
To the hyperbolic $3$-manifold with totally geodesic boundary $N^{\prime\prime}$, there is an associated complete geometrically finite hyperbolic $3$-manifold (without boundary) $N$, given by attaching a Fuchsian end to each boundary component of $N^{\prime\prime}$; in particular, $CC(N)$ is isometric to $N^{\prime\prime}$, so $CC(N)$ indeed has boundary given by a union of totally geodesic thrice punctured spheres.
Finally, let $h_{q}\in N$ denote the image of $h_{q^{\prime}}^{\prime}$ under the identifications and inclusion of $N^{\prime\prime}$ into $N$. Since $B_{N^{\prime}}(q^{\prime},R)\subseteq N_{P}^{\prime}$ we have
$B_{N^{\prime}}(q^{\prime},R)\subseteq N_{P}^{\prime}\hookrightarrow N^{\prime\prime}\hookrightarrow N$
where each of the inclusions is by isometry, so Lemma 2.3 tells us that $(N,h_{q})$ is $(\varepsilon,R)$-close to $(M,p)$, since $(N^{\prime},h_{q^{\prime}}^{\prime})$ is. ∎
Let $\mathcal{P}$ be the set of isometry classes of finite volume hyperbolic $3$-manifolds with non-empty boundary which consists of a disjoint union of totally geodesic thrice punctured spheres. Equivalently, $\mathcal{P}$ is the collection of convex cores of geometrically finite hyperbolic $3$-manifolds $N$ such that $\partial CC(N)$ is a disjoint union of thrice punctured spheres.
Lemma 3.9.
$\mathcal{P}$ is countable.
Proof.
Define the “doubling” map
[…]
as follows: for $M\in\mathcal{P}$, construct $D(M)$ by taking a copy, $\overline{M}$, of $M$ with the reversed orientation, and gluing $M$ to $\overline{M}$ by the identity map between the boundaries. Recall that the codomain of $D$ is countable.
Choose $M\in\mathcal{P}$. The doubling process determines a set of finitely many embedded totally geodesic thrice punctured spheres in $D(M)$. Each such thrice punctured sphere corresponds to a conjugacy class of $2$-generated subgroups of $\pi_{1}(D(M))$, of which there are only countably many (there are in fact only finitely many, though we will not need this fact). The map $D$ is therefore countable-to-$1$, so $\mathcal{P}$ is countable. ∎
We can now complete the construction of our dense leaf.
Theorem B.
There exists a hyperbolic $3$-manifold $M$ such that the leaf $\ell(M)$ is dense in $\mathcal{H}_{\infty}$.
Proof.
By Lemma 3.9, we may enumerate $\mathcal{P}=\{M_{0},M_{1},M_{2},...\}$. Note that each $M_{i}\in\mathcal{P}$ has an even number of boundary components, since the punctures of the thrice punctured sphere boundary components pair up into rank-1 cusps in $M_{i}$. In particular, for each $i\geq 0$ we can select distinct components $L_{i}$ and $R_{i}$ of $\partial M_{i}$. All totally geodesic thrice punctured spheres are isometric, so we may construct a new manifold $M^{\prime}$ by gluing $M_{i}$ to $M_{i+1}$ by an orientation reversing isometry taking $R_{i}$ to $L_{i+1}$ for all $i\geq 0$. See Figure 2 for a depiction of the gluing.
Figure 2: Some possible structures for the $\partial M_{i}$’s are depicted, along with the first few gluings in the construction of $M^{\prime}$.
Observe that $M^{\prime}$ is a hyperbolic $3$-manifold with boundary given by a union of thrice punctured spheres. We now construct $M$ by gluing a Fuchsian end with totally geodesic thrice punctured sphere boundary to each component of the boundary of $M^{\prime}$, again via an orientation reversing isometry from boundary to boundary. Once we check the completeness of $M$, it will follow that $M$ is the hyperbolic $3$-manifold such that $CC(M)$ is isometric to $M^{\prime}$. Note that $M$ is orientable.
To check the completeness of $M$, it will suffice to show that the cusp neighborhoods in $M$ are foliated by Euclidean tori or annuli (see, for example, [THUR, Section 3.10]). For each piece in the gluing, either one of the $M_{i}$ or a Fuchsian end, any cusp in that piece is foliated by Euclidean tori or annuli. Any such annulus $S^{1}\times I$ intersects the boundary of that piece orthogonally, and the annulus is uniquely defined within that cusp neighborhood by its “circumference,” the minimal Euclidean length in the $S^{1}$ direction. In each gluing, each annulus is glued to another annulus with the same circumference, hence these annuli piece together to form Euclidean annuli or tori, which confirms that $M$ is complete.
Let us now check that $\ell(M)$ is dense in $\mathcal{H}_{\infty}$. Choose $\varepsilon,R>0$ and $(N^{\prime},h^{\prime})\in\mathcal{H}_{\infty}$. Proposition 3.8 tells us that there is some $(N,h_{q})\in\mathcal{H}_{\infty}$ such that $CC(N)\in\mathcal{P}$, $B_{N}(q,R)\subseteq CC(N)$, and $(N,h_{q})$ is $(\varepsilon,R)$-close to $(N^{\prime},h^{\prime})$. By construction, $CC(N)$ embeds isometrically into $M$. Let $f\in\mathcal{F}M$ be the image of $h_{q}$ under this isometric embedding. Then, Lemma 2.3 tells us that $(M,f)$ is $(\varepsilon,R)$-close to $(N^{\prime},h^{\prime})$, since $(N,h_{q})$ is. Hence, we have shown that $\ell(M)$ is a dense leaf of $\mathcal{H}_{\infty}$. ∎
Note that by permuting the enumeration of $\mathcal{P}$ in the proof of Theorem B, one produces uncountably many distinct leaves which are dense in $\mathcal{H}_{\infty}$.
4 Paths
Here we will study paths in $\mathcal{H}$, starting with Section 4.1 in which we will present a number of examples. In addition to general interest, these examples will motivate some of our theorems and demonstrate sharpness of some hypotheses. Section 4.2 will develop the general theory that we will use to understand constraints on possible paths in $\mathcal{H}$. In Section 4.3 we will analyze paths of convex cocompact Kleinian groups (Theorem C), and Section 4.4 will conclude with constructing an infinite family of path components of $\mathcal{H}_{\infty}$ (Theorem D), in particular demonstrating that $\mathcal{H}_{\infty}$ is not path connected (Corollary E).
4.1 Examples
We have already come across the first example, but we will briefly record it here.
Example A.1: (Moving frame) For a fixed hyperbolic $3$-manifold $M$ and path $\gamma:[0,1]\rightarrow\mathcal{F}M$, the map $t\mapsto(M,\gamma(t))$ is a path in $\mathcal{H}$. The fact that this describes a path follows directly from Proposition 3.1, which establishes the continuity of the map $L_{M}:\mathcal{F}M\rightarrow\mathcal{H}$. Since a change of baseframe corresponds to conjugating the corresponding Kleinian group, this example corresponds to fixing a Kleinian group $\Gamma$ and a path $t\mapsto\psi_{t}\in\text{PSL}_{2}\mathbb{C}$ and taking the path in the geometric topology on $\mathcal{D}$ given by $t\mapsto\psi_{t}\Gamma\psi_{t}^{-1}$.
Example A.2: (Frame to infinity) Suppose that for a fixed hyperbolic $3$-manifold $M$ and smooth path $\gamma:[0,\infty)\rightarrow M$, the injectivity radius of $M$ along $\gamma$ satisfies $\text{inj}_{M}(\gamma(t))\rightarrow\infty$ as $t\rightarrow\infty$. Then, as in Lemma 3.3, if $f\in F_{\gamma(0)}M$ and $f^{t}$ is the time $t$ parallel transport of $f$ along $\gamma$, then $(M,f^{t})$ is a path in $\mathcal{H}$ which converges to $(\mathbb{H}^{3},\mathcal{O}_{O})$.
Example A.2 is an initial example of the topology of the underlying manifold changing along a path in $\mathcal{H}$. As seen in Lemma 3.3, a path such as $\gamma$ may be constructed in $M=\mathbb{H}^{3}/\Gamma$ for any Kleinian group $\Gamma$ with $\Omega(\Gamma)\neq\varnothing$.
Example B.1: (Quasi-conformal deformations) Suppose $\Gamma$ is a nonelementary geometrically finite Kleinian group. For any path of geometrically finite representations $G:[0,1]\rightarrow GF(\Gamma)$ (see Section 2.3 for definitions), the map $t\mapsto G(t)(\Gamma)$ is a path in $\mathcal{D}$.
Define $G_{im}:[0,1]\rightarrow\mathcal{D}$ by $G_{im}(t)=(G(t))(\Gamma)$. The proof of Marden’s stability theorem [MAR] demonstrates that the Dirichlet fundamental polyhedra with respect to the fixed basepoint $O\in\mathbb{H}^{3}$ for $G_{im}(t)$ vary uniformly continuously on compact balls about $O\in\mathbb{H}^{3}$ in the Hausdorff topology. In particular, $G_{im}(t)$ is a path of Kleinian groups with respect to the topology of polyhedral convergence (see [MAT]), hence is a path of Kleinian groups in the geometric topology on $\mathcal{D}$. In other words, $G$ may also be viewed as a path into $S(\Gamma)$ equipped with the topology of strong convergence. We also note that the continuity of $G_{im}$ with respect to the geometric topology follows immediately from a result of Anderson and Canary on strong convergence [ANCAN], since each $G(t)$ is type preserving. Following from the correspondence between quasiconformal deformations and quasi-isometries [RENORM], the path in $\mathcal{H}$ corresponding to $G_{im}$ is a path of quasi-isometric framed hyperbolic $3$-manifolds, where cusps map to cusps.
Example B.2: (Cusp rank) Let $\Gamma$ be a geometrically finite Kleinian group with a parabolic element $\psi\in\Gamma$ such that $\mathbb{H}^{3}/\Gamma$ has a rank-$1$ cusp corresponding to $\psi$; more precisely, $\psi$ generates the largest abelian subgroup of $\Gamma$ containing $\psi$. Then, there exists a path $\varphi:[t_{0},\infty)\rightarrow\text{PSL}_{2}\mathbb{C}$ of parabolic isometries so that $\langle\psi,\varphi(t)\rangle$ is a rank-$2$ abelian group for all $t$ and $t\mapsto\langle\Gamma,\varphi(t)\rangle$ is a path in $\mathcal{D}$ converging to $\Gamma$ as $t\rightarrow\infty$.
This example is discussed in detail by Fuchs, Purcell and Stewart in [PUR, Section 4], and we will simply expand on some details of their analysis, with a focus on the topology of the approximating manifolds. By conjugating $\Gamma$, we may assume $\psi$ corresponds to the Möbius transformation $z\mapsto z+1$. For $t>0$ let $\varphi_{t}$ denote the element of $\text{PSL}_{2}\mathbb{C}$ corresponding to $z\mapsto z+it$. Using the upper half space model for $\mathbb{H}^{3}$, define the subsets
$B_{\pm}=B_{\pm}(t)=\{(z,h)\in\mathbb{C}\times\mathbb{R}_{+}=\mathbb{H}^{3}\;|\;\pm\text{Im}(z)\geq t/2\}.$
An analysis (see [PUR]) of the standard cusp regions corresponding to $\psi$ as defined in [BOWGF] tells us that for $t$ sufficiently large $B_{+},B_{-}$, and $B_{+}\cup B_{-}$ are each precisely invariant for $\langle\psi\rangle$ in $\Gamma$. For such $t$, since $B_{+}=\overline{\mathbb{H}^{3}\backslash\varphi_{t}(B_{-})}$, the second Klein-Maskit combination theorem (see [MAT], [MASK], [ABI]) tells us that the group $\Gamma_{t}=\langle\Gamma,\varphi_{t}\rangle$ is discrete. Fix $t_{0}>0$ so that $t$ is “sufficiently large” for all $t\geq t_{0}$: $\Gamma_{t}\in\mathcal{D}$ for all $t\geq t_{0}$. That $t\mapsto\Gamma_{t}$ describes a path in $\mathcal{D}$ for $t\in[t_{0},\infty)$ follows from Example B.1.
Before we prove that $\Gamma_{t}\rightarrow\Gamma$ as $t\rightarrow\infty$, let us discuss how to construct $M_{t}=\mathbb{H}^{3}/\Gamma_{t}$ from $M=\mathbb{H}^{3}/\Gamma$. Let $\pi:\mathbb{H}^{3}\rightarrow M$ denote the projection map corresponding to $\Gamma$. One constructs $M_{t}$ by gluing the two boundary components of $M\backslash(\pi(B_{+})\cup\pi(B_{-}))$. In the language of pared manifolds (see [THUI, Section 7]), $M$ is homeomorphic to the interior of a pared closed $3$-manifold $(\overline{M},P)$, and the rank-$1$ cusp associated to $\psi$ corresponds to a component of $P$ which is incompressible annulus $A$ contained in a surface $S\subseteq\partial\overline{M}$. Let $c\subseteq S$ denote a curve representing the homotopy class of a core curve of $A$. The surface $S\subseteq\partial\overline{M}$ has a neighborhood in $\overline{M}$ parameterized by $S\times[0,1]$, where $S=S\times\{1\}\subseteq\partial\overline{M}$. Let $\overline{M}^{\prime}$ denote the closed $3$-manifold given by drilling out an open tubular neighborhood of $c\times\{1/2\}$; assume this neighborhood is chosen to be sufficiently small to be contained in $S\times[1/4,3/4]$, and let $T$ be the toroidal boundary of this neighborhood. Set $P^{\prime}=(P\backslash A)\cup T\subseteq\partial\overline{M}^{\prime}$. A cartoon of the local picture of the construction of $(\overline{M}^{\prime},P^{\prime})$ from $(\overline{M},P)$ is given in Figure 3. Then, the pared closed $3$-manifold corresponding to each $M_{t}$ is $(\overline{M}^{\prime},P^{\prime})$.
Figure 3: This depicts the construction of $(\overline{M}^{\prime},P^{\prime})$ from $(\overline{M},P)$ in Example 2.2, depicting $S$ as a genus-$2$ surface and the paring as shaded.
Let us confirm that $\Gamma_{t}\rightarrow\Gamma$ as $t\rightarrow\infty$. Note that for large enough $t$, $O\not\in B_{+}(t)\cup B_{-}(t)$, where ${\mathcal{O}}_{O}$ is our fixed baseframing for $\mathbb{H}^{3}$. From the above construction of $M_{t}$ from $M$, we can then see that for any $R>0$, if $t$ is sufficiently large, the inclusion map $B_{\mathbb{H}^{3}}(O,R)\hookrightarrow\mathbb{H}^{3}$ descends to an embedding $B_{M}(\pi(O),R)\hookrightarrow M_{t}$. Thus, Definition 2.1 directly tells us that $\Gamma_{t}\rightarrow\Gamma$ as $t\rightarrow\infty$.
Example B.3: (Non-tame limit) For a closed surface $S$ of genus at least two, let $C\subseteq S\times\mathbb{Z}\subseteq S\times\mathbb{R}$ be a collection of disjoint simple closed curves so that no two curves in $C$ are isotopic in $S\times\mathbb{R}$ and the curves in $C_{n}=C\cap(S\times\{n\})$ fill the surface $S_{n}=S\times\{n\}$ for all $n\in\mathbb{Z}$. Then, there exists a path $G:[1,\infty]\rightarrow\mathcal{D}$ so that $G(t)$ is geometrically finite for all $t<\infty$ and $\mathbb{H}^{3}/G(\infty)$ is isometric to the unique hyperbolic structure on $(S\times\mathbb{R})\backslash C$.
For the construction, it will be slightly easier to replace the assumption that $C_{0}$ fills $S_{0}$ with the assumption that $C_{0}=\varnothing$; this simply amounts to a reparameterization of $S\times\mathbb{R}$ and $C$, so nothing is lost. For each $n$, let $P_{n}\subseteq S_{n}$ denote a tubular neighborhood of $C_{n}$ as a submanifold of $S_{n}$, so that $P_{n}$ is a disjoint union of annuli whose core curves are the elements of $C_{n}$. It follows, for example by [THUII, Theorem 6.4], that there is a hyperbolic structure, call it $M_{1}$, on the interior of the pared compact $3$-manifold $(S\times[-1,1],P_{-1}\cup P_{1})$, so that the annuli in the paring correspond to rank-$1$ cusps in $M_{1}$; note that since $C_{-1}$ and $C_{1}$ fill $S_{-1}$ and $S_{1}$ respectively, $\partial CC(M_{1})$ must be a disjoint union of totally geodesic thrice punctured spheres, and hence the hyperbolic structure on $M_{1}$ is uniquely determined. Fix a Kleinian group $\Gamma_{1}\in\mathcal{D}$ so that $M_{1}=\mathbb{H}^{3}/\Gamma_{1}$.
Let $T\subseteq S\times\mathbb{R}$ be a small tubular neighborhood of $C$ so that $\partial T$ is a disjoint union of tori and $(S\times\mathbb{R})\backslash C$ is homeomorphic to the interior of $(S\times\mathbb{R})\backslash T$. Additionally, let $T_{n}$ denote the union of the components of $T$ corresponding to curves in $C_{n}$, and assume $T_{n}\subseteq S\times[n-(1/4),n+(1/4)].$
For $n\geq 1$, assume that we have inductively constructed a Kleinian group $\Gamma_{n}$ so that $\Gamma_{n-1}\leq\Gamma_{n}$ (where $\Gamma_{0}=\{\mathds{1}\}$) and $M_{n}=\mathbb{H}^{3}/\Gamma_{n}$ gives a hyperbolic structure on interior of the pared $3$-manifold
$\overline{M_{n}}=(S\times[-n,n]\backslash\textstyle\bigcup_{-n<j<n}T_{j},P_{-n}\cup P_{n}\cup(\textstyle\bigcup_{-n<j<n}\partial T_{j})),$
relative to the paring. The hyperbolic structure on $M_{n}$ must be unique, since $\partial CC(M_{n})$ is a disjoint union of totally geodesic thrice punctured spheres. For each rank-$1$ cusp of $M_{n}$, Example B.2 allows us to bring in a parabolic isometry from infinity to construct a path in $\mathcal{D}$ from $\Gamma_{n}$ to $\Gamma_{n}^{\prime}$ so that $\Gamma_{n}$ is contained in all intermediate groups, including $\Gamma_{n}^{\prime}$, and $M_{n}^{\prime}=\mathbb{H}^{3}/\Gamma_{n}^{\prime}$ is homeomorphic to the interior of $S\times[-(n+1),n+1]\backslash(\bigcup_{-n\leq j\leq n}T_{n})$. Note that $\Gamma_{n}^{\prime}$ is not uniquely defined: we bring in each new parabolic isometry along a path from infinity in $\text{PSL}_{2}\mathbb{C}$, but there is choice in how “far” to bring each in along their respective paths. Note, however, that the rigidity of $M_{n}$ implies that $CC(M_{n})$ embeds isometrically into $M_{n}^{\prime}$ under the natural quotient map $M_{n}\rightarrow M_{n}^{\prime}$.
By work of Ahlfors, Bers, Marden, and Sullivan (see eg. [MAT]) we may identify $\overline{GF}(\Gamma_{n}^{\prime})$ with the Teichmüller space $\mathcal{T}(S_{-(n+1)})\times\mathcal{T}(S_{n+1})$. The pants decompositions of $S_{-(n+1)}$ and $S_{n+1}$ given by $C_{-(n+1)}$ and $C_{n+1}$ respectively provide Fenchel-Nielsen coordinates [BEN] for the respective Teichmüller spaces. Consider a path $t\mapsto[\rho_{t}]\in\overline{GF}(\Gamma_{n}^{\prime})$ corresponding to sending the lengths of the curves in $C_{-(n+1)}$ and $C_{n+1}$ to zero as $t\rightarrow\infty$, and, say, keeping the twist parameters constant. There is a unique lift to a path $t\mapsto\rho_{t}\in GF(\Gamma_{n}^{\prime})$ so that $\rho_{t}|_{\Gamma_{n}}=\mathds{1}_{\Gamma_{n}}$ for all $t$, by Lemma 2.7 and the rigidity of $CC(M_{n})$. By Example B.1, the map $t\mapsto\rho_{t}(\Gamma_{n}^{\prime})\in\mathcal{D}$ is a path in $\mathcal{D}$.
Observe that the hypothesis that all curves in $C$ are non-isotopic guarantees that as a pared $3$-manifold, $\overline{M_{n+1}}$ is acylindrical and has incompressible boundary. In the language of [EGG], this implies that the gallimaufry associated to the paring on $\overline{M_{n+1}}$ is doubly incompressible. In particular, it then follows from [EGG] and Lemma 2.7 that as $t\rightarrow\infty$, the path $t\mapsto\rho_{t}(\Gamma_{n}^{\prime})$ converges to a Kleinian group $\Gamma_{n+1}$ so that $M_{n+1}=\mathbb{H}^{3}/\Gamma_{n+1}$ gives a hyperbolic structure on the interior of the pared hyperbolic $3$-manifold $\overline{M_{n+1}}$. Noting that $\Gamma_{n}\leq\Gamma_{n+1}$, the inductive step is complete.
Hence, we may concatenate and reparameterize the above paths to define a path $G:[1,\infty)\rightarrow\mathcal{D}$ so that $G(n)=\Gamma_{n}$ for all $n$ and $\Gamma_{n}\leq G(t)$ for all $t\geq n$. Additionally, observe that for all $t\geq n$ the natural projection $M_{n}\rightarrow\mathbb{H}^{3}/G(t)$ embeds $CC(M_{n})$ isometrically into $\mathbb{H}^{3}/G(t)$. By Lemma 2.3, the path $G(t)$ converges to some $\Gamma_{\infty}$ where $M_{\infty}=\mathbb{H}^{3}/\Gamma_{\infty}$ is homeomorphic to $(S\times\mathbb{R})\backslash C$.
An example of a collection of curves such as those in Example B.3 may be constructed by picking a pants decomposition $D\subseteq S$ of $S$, a pseudo-Anosov homeomorphism $f:S\rightarrow S$, and letting $C=\bigcup(f^{n}(D)\times\{n\})$. In this case, the hyperbolic structure on $M_{\infty}$, which is homeomorphic to $(S\times\mathbb{R})\backslash C$, has a $\mathbb{Z}$-symmetry. In particular, each piece $(S\times[n,n+1])\backslash C$ is isometric.
We also note that the manifolds $M_{\infty}$ (constructed from general $C$, not just via a pseudo-Anosov) are among those non-tame hyperbolic $3$-manifolds which Thurston approximates with sequences of quasi-Fuchsian manifolds in [THUII, Section 7], though the techniques and goals are quite different.
Example C.1: (Schottky) In this example, we will construct a path of rank-2 Schottky (free) groups that converges to an infinite cyclic group. For $s>1$, define
$\psi_{s}=\begin{bmatrix}s&0\\ s&1/s\end{bmatrix}\in\text{PSL}_{2}\mathbb{R}.$
As discussed in [FORD], if
$C_{s}=\{z\in\mathbb{C}\;|\;|z+1/s^{2}|=1/s\}\;\;\text{ and }\;\;C_{s}^{\prime}=\{z\in\mathbb{C}\;|\;|z-1|=1/s\},$
then $\psi_{s}$ acts on the Riemann sphere $S^{2}_{\infty}$ by mapping the exterior of $C_{s}$ onto the interior of $C_{s}^{\prime}$.
Let $D$ and $D^{\prime}$ be two circles on the Riemann sphere whose designated interiors are disjoint from each other and disjoint from $\{z\in\mathbb{C}:|z|\leq 2\}$ (and hence disjoint from $C_{s},C_{s}^{\prime}$ for all $s>1$). Now, choose $\varphi\in\text{PSL}_{2}\mathbb{C}$ to be any isometry taking the exterior of $D$ to the interior of $D^{\prime}$.
For $t\in(1,\infty]$, let $\Gamma_{t}\leq\text{PSL}_{2}\mathbb{C}$ be the subgroup given by
$\Gamma_{t}=\begin{cases}\langle\varphi,\psi_{t}\rangle,&1<t<\infty\\ \langle\varphi\rangle,&t=\infty.\end{cases}$
It is classical, following from the first Klein-Maskit combination theorem ([MAT], [MASK], [ABI]), that each $\Gamma_{t}$ is a discrete free group. That $t\mapsto\Gamma_{t}$ is a path in $\mathcal{D}$ on the domain $t\in(1,\infty)$ follows from Example B.1. To see that $\Gamma_{t}\rightarrow\Gamma_{\infty}$ as $t\rightarrow\infty$, note that each word $w$ in the abstract free group $\langle a,b\rangle$ corresponds to a path $t\mapsto w_{t}$ in $\text{PSL}_{2}\mathbb{C}$ with domain $t\in(0,\infty)$ under the isomorphism $\langle a,b\rangle\rightarrow\Gamma_{t}$ given by $a\mapsto\varphi$, $b\mapsto\psi_{t}$. For any reduced word $w\in\langle a,b\rangle\backslash\langle a\rangle$, the path $w_{t}$ must diverge to $\infty$ as $t\rightarrow\infty$; for example, note that a loop in $\mathbb{H}^{3}/\Gamma_{t}$ representing the conjugacy class of $w_{t}\in\Gamma_{t}\cong\pi_{1}(\mathbb{H}^{3}/\Gamma_{t})$ must have length at least that of the closed geodesic corresponding to $\psi_{t}$. Hence, the only elements of $\text{PSL}_{2}\mathbb{C}$ that may be accumulated on by the $\Gamma_{t}$’s as $t\rightarrow\infty$ are the elements of $\langle\varphi\rangle$, so it follows from Fact 2.4 that $\Gamma_{t}\rightarrow\Gamma_{\infty}$ as $t\rightarrow\infty$.
Example C.2: (Convex cocompact free product) For Kleinian groups $G$ and $H$ with $\Omega(G),\Omega(H)\neq\varnothing$, there exists a path $\varphi:[0,\infty)\rightarrow\text{PSL}_{2}\mathbb{C}$ so that if $H_{t}=\varphi(t)H\varphi(t)^{-1}$, then the following hold:
  1. 1.
    each $\langle G,H_{t}\rangle$ is discrete and splits as a free product $\langle G,H_{t}\rangle=G*H_{t}$,
  2. 2.
    $t\mapsto G*H_{t}$ is a path in $\mathcal{D}$,
  3. 3.
    $G*H_{t}\rightarrow G$ in $\mathcal{D}$ as $t\rightarrow\infty$.
We may choose an open disk $D_{1}\subseteq\Omega(G)$ such that $D_{1}$ is precisely invariant under $\{\mathds{1}\}$ in $G$. Fix $x_{1},x_{2}\in D_{1}$. By replacing $H$ with a $\text{PSL}_{2}\mathbb{C}$-conjugate subgroup, we may assume that $x_{1}\in\Omega(H)$. Now, we may choose an open disk $D_{2}\subseteq\Omega(H)$ centered at $x_{1}$ such that $D_{2}$ is precisely invariant under $\{\mathds{1}\}$ in $H$.
For $t>0$, let $\varphi_{t}$ be the hyperbolic isometry with repelling fixed point $x_{1}$, attracting fixed point $x_{2}$, no torsion part, and hyperbolic translation length $t$ along the geodesic limiting to $x_{1}$ and $x_{2}$. For $t>0$, define
$\Gamma_{t}=\langle G,\varphi_{t}H\varphi_{t}^{-1}\rangle.$
For $t$ sufficiently large, say for $t\geq t^{*}$, we have $S^{2}_{\infty}\backslash\overline{D_{1}}\subseteq\varphi_{t}(D_{2})$. Since $D_{2}$ is precisely invariant under $\{\mathds{1}\}$ in $H$, it follows that for $t\geq t^{*}$, $S^{2}_{\infty}\backslash\overline{D_{1}}$ is precisely invariant under $\{\mathds{1}\}$ in $\varphi_{t}H\varphi_{t}^{-1}$. Hence, since $D_{1}$ is precisely invariant under $\{\mathds{1}\}$ in $G$, the first Klein-Maskit combination theorem, as stated in [ABI], implies that $\Gamma_{t}$ is discrete for $t\geq t^{*}$, with $\Gamma_{t}\cong G*\varphi_{t}H\varphi_{t}^{-1}\cong G*H$. A similar argument as made in Example C.1 shows that $\Gamma_{t}$ is a path in $\mathcal{D}$ on the domain $t\in[t^{*},\infty)$ and that $\Gamma_{t}\rightarrow G$ as $t\rightarrow\infty$.
Let us also note what is occurring in the corresponding path of framed manifolds in $\mathcal{H}$. Let $N_{1}=\mathbb{H}^{3}/G$, $N_{2}=\mathbb{H}^{3}/H$, and $M_{t}=\mathbb{H}^{3}/\Gamma_{t}$, for $t\geq t^{*}$. The convex cores $CC(N_{1})$ and $CC(N_{2})$ isometrically embed into $M_{t}$ for all $t$, with their images getting farther apart as $t\rightarrow\infty$. In particular, if $f^{t}\in\mathcal{F}M_{t}$ is the $\Gamma_{t}$-projection of $\mathcal{O}_{O}$, then the embedded $CC(N_{1})\hookrightarrow M_{t}$ stays fixed relative to $f^{t}$, while the embedded $CC(N_{2})\hookrightarrow M_{t}$ gets infinitely far away from $f^{t}$, disappearing and leaving behind $N_{1}$ as $t\rightarrow\infty$.
4.2 Path Machinery
This subsection will develop some machinery for discussing paths in $\mathcal{H}$ and will establish some preliminary results on how these paths behave. First, we will have a lemma concerning sequences in $\mathcal{D}$.
Lemma 4.1.
Fix $\Gamma\in\mathcal{D}$. For each $\psi\in\Gamma$ there exists a neighborhood $U_{\psi}\subseteq\text{PSL}_{2}\mathbb{C}$ such that if $\Gamma_{n}\rightarrow\Gamma$ is a convergent sequence in $\mathcal{D}$, then $|\Gamma_{n}\cap U_{\psi}|=1$ for $n$ sufficiently large.
Proof.
[…]
Now, suppose for contradiction that $\{\Gamma_{n}\}$ is a sequence in $\mathcal{D}$ converging to $\Gamma$ such that there exists $\psi_{n}\in(\Gamma_{n}\cap U)\backslash\{\mathds{1}\}$ for all $n$. For each $n$, define $v_{n}\in B_{r/2}$ such that $\text{exp}(v_{n})=\psi_{n}$, and choose $m_{n}\in\mathbb{Z}_{+}$ such that $r/2\leq\|m_{n}\cdot v_{n}\|<r$. Then,
$\psi_{n}^{m_{n}}=\text{exp}(m_{n}\cdot v_{n})\in\overline{U^{\prime}\backslash U}\cap\Gamma_{n}.$
Since $\overline{U^{\prime}\backslash U}$ is compact, the sequence $\{\psi_{n}^{m_{n}}\}_{n\geq 1}$ has an accumulation point $\psi^{*}$ in $\overline{U^{\prime}\backslash U}$, so the first condition for convergence in the geometric topology in Fact 2.4 implies that $\psi^{*}\in\overline{U^{\prime}\backslash U}\cap\Gamma$. This, however, is a contradiction, so taking $U_{\mathds{1}}=U$ suffices.
To address case of general $\psi\in\Gamma$, first consider the map $F:\text{PSL}_{2}\mathbb{C}\times\text{PSL}_{2}\mathbb{C}\rightarrow\text{PSL}_{2}\mathbb{C}$ given by $F(\varphi,\tau)=\varphi^{-1}\tau$. Note that $F$ is continuous, so $F^{-1}(U)$ is an open neighborhood of $\mathds{1}\times\mathds{1}$, where $U$ is the same subset as in the previous paragraphs. We can then find $U^{*}\subseteq U$ containing $\mathds{1}$ such that $U^{*}\times U^{*}\subseteq F^{-1}(U)$. Observe then that $\varphi^{-1}\tau\in U$ for all $\varphi,\tau\in U^{*}$.
For $\psi\in\Gamma\backslash\{\mathds{1}\}$, set $U_{\psi}=\psi\cdot U^{*}$. As above, consider a sequence $\{\Gamma_{n}\}$ converging in $\mathcal{D}$ to $\Gamma$. If $\varphi_{n},\tau_{n}\in\Gamma_{n}\cap U_{\psi}$ for some $n$, then $\psi^{-1}\varphi_{n},\psi^{-1}\tau_{n}\in U^{*}$, so the choice of $U^{*}$ tells us
$\varphi_{n}^{-1}\tau_{n}=(\psi^{-1}\varphi_{n})^{-1}(\psi^{-1}\tau_{n})\in U=U_{\mathds{1}}.$
We know from above that for $n$ sufficiently large, we must have $\Gamma_{n}\cap U=\{\mathds{1}\}$, so it follows that $|\Gamma_{n}\cap U_{\psi}|\leq 1$ for $n$ sufficiently large. Since $\Gamma_{n}\rightarrow\Gamma$, the second condition for convergence in the geometric topology given in Fact 2.4 tells us that we in fact have $|\Gamma_{n}\cap U_{\psi}|=1$ for $n$ sufficiently large. ∎
Definition 4.2.
Let $\Gamma:[a,b]\rightarrow\mathcal{D}$ be a path. For $s\in[a,b]$ and $\psi\in\Gamma(s)$, we will call a path $j:I\rightarrow\text{PSL}_{2}\mathbb{C}$ a $\Gamma$-path through $\psi$ based at $s$ if $j$ satisfies the following conditions:
  1. i.
    $I\subseteq[a,b]$ is an interval containing $s$
  2. ii.
    $j(t)\in\Gamma(t)$ for all $t\in I$
  3. iii.
    $j(s)=\psi$.
For the remainder of the subsection, fix a path $\Gamma:[0,1]\rightarrow\mathcal{D}$, and let $\Gamma_{t}=\Gamma(t)$. The following proposition gives local existence of $\Gamma$-paths through $\psi$ based at $s$ for any $s\in[0,1]$ and $\psi\in\Gamma_{s}$. This may be interpreted as a path analogue of the second condition for sequential convergence in the geometric topology on $\mathcal{D}$, as stated in Fact 2.4.
Proposition 4.3.
For all $s\in[0,1]$ and $\psi\in\Gamma_{s}$, there exists an interval open in $[0,1]$ which is the domain of a $\Gamma$-path through $\psi$ based at $s$. Further, if $j:I\rightarrow\text{PSL}_{2}\mathbb{C}$ is any $\Gamma$-path through $\psi$ based at $s$, $j$ is the unique $\Gamma$-path through $\psi$ based at $s$ with domain $I$.
Proof.
First, fix a neighborhood $U_{\psi}\subseteq\text{PSL}_{2}\mathbb{C}$ of $\psi$ satisfying the conclusion of Lemma 4.1, with respect to sequences converging to $\Gamma_{s}$. Now, observe that for $t$ close enough to $s$, we have $|\Gamma_{t}\cap U_{\psi}|=1$. Hence, for some open interval $I^{\prime}$ containing $s$, we may define a function $j:I^{\prime}\rightarrow\text{PSL}_{2}\mathbb{C}$ by setting $j(t)$ to be the unique element in $\Gamma_{t}\cap U_{\psi}$. In particular, $j(s)=\psi$. The fact that $j$ is continuous follows from $\Gamma$ being a path in $\mathcal{D}$. Hence, $j$ is a $\Gamma$-path through $\psi$ based at $s$, and observe that $j$ is the uniquely defined such path on this domain.
Now, fix an interval $I\subseteq[0,1]$ containing $s$ and suppose $j_{1},j_{2}:I\rightarrow\text{PSL}_{2}\mathbb{C}$ are $\Gamma$-paths through $\psi$ based at $s$. Let $t_{*}=\inf\{t\in I\;|\;j_{1}(t^{\prime})=j_{2}(t^{\prime})\text{ for all $t^{\prime}\in[t,s]$}\}$. If $t_{*}\neq\inf I$, then continuity implies that $j_{1}(t_{*})=j_{2}(t_{*})$ and it follows that $j_{1}$ and $j_{2}$ are $\Gamma$-paths through $j_{1}(t_{*})$ based at $t_{*}$. The preceding paragraph then implies that $j_{1}$ and $j_{2}$ must agree near $t_{*}$, contradicting the definition of $t_{*}$. Hence, $t_{*}=\inf I$, and, with an analogous argument with a supremum, the uniqueness claim follows. ∎
Notice that Proposition 4.3 implies that for each $s\in[0,1]$ and $\psi\in\Gamma_{s}$, there is an interval $I_{\psi}^{s}\subseteq[0,1]$ which is the maximal domain on which we can define a $\Gamma$-path through $\psi$ based at $s$, call it $j_{\psi}^{s}:I_{\psi}^{s}\rightarrow\text{PSL}_{2}\mathbb{C}$. For an interval $I\subseteq[0,1]$, we define the frontier $\text{Fr}(I)=\overline{I}\backslash\text{Int}_{[0,1]}(I)$, where $\text{Int}_{[0,1]}(I)$ is the interior of $I$ with respect to the topology on $[0,1]$. Note that $\text{Fr}(I)\subseteq\{\inf I,\sup I\}$.
Lemma 4.4.
Fix $s\in[0,1]$ and $\psi\in\Gamma_{s}$.
  1. 1.
    The interval $I_{\psi}^{s}$ is open in the topology on $[0,1]$. In particular, if $\inf I_{\psi}^{s}\not\in\text{Fr}(I_{\psi}^{s})$, then $\inf I_{\psi}^{s}=0$; similarly, if $\sup I_{\psi}^{s}\not\in\text{Fr}(I_{\psi}^{s})$, then $\sup I_{\psi}^{s}=1$.
  2. 2.
    As $t\in I_{\psi}^{s}$ converges to a point in $\text{Fr}(I_{\psi}^{s})$, $j_{\psi}^{s}(t)$ diverges to $\infty$ in $\text{PSL}_{2}\mathbb{C}$.
Proof.
The first claim follows immediately from Proposition 4.3, which tells us that $I_{\psi}^{s}$ is a union of intervals containing $s$ that are open in $[0,1]$.
For the second claim, fix $t^{\prime}\in\text{Fr}(I_{\psi}^{s})$ and suppose $j_{\psi}^{s}(t)$ does not diverge to $\infty$ as $t\in I_{\psi}^{s}$ converges to $t^{\prime}$. Then, as $t$ converges to $t^{\prime}$, $j_{\psi}^{s}(t)$ accumulates on some $\varphi\in\text{PSL}_{2}\mathbb{C}$. By the first condition for convergence in the geometric topology on $\mathcal{D}$, we must then have $\varphi\in\Gamma_{t^{\prime}}$. Let $U_{\varphi}\subseteq\text{PSL}_{2}\mathbb{C}$ be the neighborhood of $\varphi$ given by Lemma 4.1 with respect to $\Gamma_{t^{\prime}}$, and note that it follows that $j_{\psi}^{s}(t)\in U_{\varphi}$ for $t$ sufficiently close to $t^{\prime}$. The proof of Lemma 4.1 shows that one can take the neighborhood $U_{\psi}$ to be arbitrarily small, so we see that $j_{\psi}^{s}(t)$ converges to $\varphi\in\Gamma_{t^{\prime}}$ as $t$ approaches $t^{\prime}$. Hence, the maximality of $I_{\psi}^{s}$ implies $t^{\prime}\in I_{\psi}^{s}$, which contradicts that $I_{\psi}^{s}$ is open. ∎
Let $\overline{\text{PSL}_{2}\mathbb{C}}=\text{PSL}_{2}\mathbb{C}\cup\{\infty\}$ denote the one-point compactification of $\text{PSL}_{2}\mathbb{C}$. For $s,t\in[0,1]$ we now define a map $J_{s,t}:\Gamma_{s}\rightarrow\overline{\text{PSL}_{2}\mathbb{C}}$ by
$J_{s,t}(\psi)=\begin{cases}j_{\psi}^{s}(t),&t\in I_{\psi}^{s}\\ \infty,&\text{else}.\end{cases}$ (1)
For each $s\in[0,1]$ and $\psi\in\Gamma_{s}$, the map $t\mapsto J_{s,t}(\psi)=j_{\psi}^{s}(t)$ on $I_{\psi}^{s}$ describes a path in $\text{PSL}_{2}\mathbb{C}$, which diverges to infinity at any points of $\text{Fr}(I_{\psi}^{s})$. Outside of $I_{\psi}^{s}$, $J_{s,t}(\psi)$ is constant at the $\infty\in\overline{\text{PSL}_{2}\mathbb{C}}$. We will record some facts about each map $J_{s,t}$, whose image is contained in $\Gamma_{t}\cup\infty$.
Lemma 4.5.
Fix $s\in[0,1]$.
  1. 1.
    The family of functions $J_{s,t}$ varies continuously in $t$ with respect to the compact-open topology.
  2. 2.
    If $t,t^{\prime}\in I_{\psi}^{s}$, then $(J_{t,t^{\prime}}\circ J_{s,t})(\psi)=J_{s,t^{\prime}}(\psi)$.
Proof.
For the first claim, it suffices to show that for all $\psi\in\Gamma_{s}$, $t\mapsto J_{s,t}(\psi)$ is a path in $\overline{\text{PSL}_{2}\mathbb{C}}$, since $\Gamma_{s}$ is discrete. By definition, $t\mapsto J_{s,t}(\psi)$ agrees with the path $j_{\psi}^{s}$ on $I_{\psi}^{s}$. Hence, it therefore remains to observe that Lemma 4.4 implies that as $t$ converges to a point in the frontier of $I_{\psi}^{s}$, $J_{s,t}(\psi)$ converges to $\infty$ in $\overline{\text{PSL}_{2}\mathbb{C}}$.
For the second claim, observe that as $t^{\prime\prime}$ ranges between $t$ and $t^{\prime}$, $t^{\prime\prime}\mapsto j_{\psi}^{s}(t^{\prime\prime})$ describes a $\Gamma$-path through $j_{\psi}^{s}(t)$ based at $t$, so the uniqueness of such paths given by Proposition 4.3 gives the desired claim. ∎
For a Kleinian group $\Gamma^{\prime}$, we will say that a map $J:\Gamma^{\prime}\rightarrow\overline{\text{PSL}_{2}\mathbb{C}}$ is finite valued if the image $J(\Gamma^{\prime})$ is contained in $\text{PSL}_{2}\mathbb{C}$. Understanding when the maps $J_{s,t}$ are finite valued will be an important tool for analyzing the path $t\mapsto\Gamma_{t}$. In particular, we will see that if $J_{s,t}$ is finite valued for all $t\in I\subseteq[0,1]$, then $J_{s,t}$ is an injective homomorphism for all $t\in I$, so the map $t\mapsto J_{s,t}\in S(\Gamma_{s})$ is a path in the topology of strong convergence on $S(\Gamma_{s})$.
Proposition 4.6.
For any $s,t\in[0,1]$ and subgroup $H\leq\Gamma_{s}$, the following are equivalent:
  1. 1.
    $J_{s,t}|_{H}$ is finite valued
  2. 2.
    $J_{s,t}|_{H}$ is an injective homomorphism
  3. 3.
    $t\in I_{\psi}^{s}$ for all $\psi\in H$.
Proof.
($3\Rightarrow 2$): Let $I_{*}=\bigcap_{\psi\in H}I_{\psi}^{s}$, and fix $\psi,\varphi\in H$. Notice that for $t\in I_{*}$, $t\mapsto J_{s,t}(\psi)\cdot J_{s,t}(\varphi)$ is a $\Gamma$-path through $\psi\varphi$ based at $s$, so the uniqueness statement in Proposition 4.3 implies that $J_{s,t}(\psi)\cdot J_{s,t}(\varphi)=J_{s,t}(\psi\varphi)$ for all $t\in I_{*}$. Thus, $J_{s,t}|_{H}$ is a homomorphism for all $t\in I_{*}$.
Now, fix $t\in I_{*}$ and suppose that $J_{s,t}(\psi)=\mathds{1}$ for some $\psi\in H$, and observe that the path $j_{\psi}^{s}$ is a $\Gamma$-path through $\mathds{1}$ based at $t$. The constant path at $\mathds{1}$ is the unique $\Gamma$-path through $\mathds{1}$ based at $t$, so we must have that $\psi=\mathds{1}$, and the injectivity of $J_{s,t}$ follows.
The implications ($2\Rightarrow 1$) and ($1\Rightarrow 3$) are immediate from the definitions. ∎
Informally, the previous proposition says that $J_{s,t}|_{H}$ will be an injective homomorphism if and only if as $t^{\prime}$ ranges from $s$ to $t$, the path $t^{\prime}\mapsto J_{s,t^{\prime}}(\psi)$ does not diverge to $\infty$ for any $\psi\in H$. Additionally, we see that for any distinct $\psi,\varphi\in\Gamma_{s}$, the paths $t\mapsto J_{s,t}(\psi)$ and $t\mapsto J_{s,t}(\varphi)$ stay away from each other on $I_{\psi}^{s}\cup I_{\varphi}^{s}$.
Corollary 4.7.
For any $s\in[0,1]$ and finitely generated subgroup $H\leq\Gamma_{s}$, the restriction $J_{s,t}|_{H}$ is an injective homomorphism for $t\in[0,1]$ sufficiently close to $s$.
Proof.
Let $\{\varphi_{1},...,\varphi_{k}\}$ be a finite set of generators for $H$ and set $I_{*}=\cap_{i=1}^{k}I_{\varphi_{i}}^{s}$. Lemma 4.4(1) tells us each $I_{\varphi_{i}}^{s}$ is open in $[0,1]$ and contains $s$, so $I_{*}$ must be open in $[0,1]$ and contain $s$. Note that if $\psi_{1},\psi_{2}\in\Gamma_{s}$, then $I_{\psi_{1}}^{s}\cap I_{\psi_{2}}^{s}\subseteq I_{\psi_{1}\psi_{2}}^{s}$, so $I_{*}\subseteq I_{\psi}^{s}$ for all $\psi\in H$. Thus, the claim follows from Proposition 4.6($3\Rightarrow 2$). ∎
We will now briefly contrast this result, Corollary 4.7, on paths in $\mathcal{D}$ with the analogous situation for convergent sequences in $\mathcal{D}$, focusing on injectivity. A classical example from Jørgensen and Marden [JORMAR] of conformal Dehn surgery constructs a sequence of infinite cyclic groups $\langle\psi_{i}\rangle$ generated by a hyperbolic converging geometrically in $\mathcal{D}$ to a rank-$2$ abelian parabolic group $\langle a,b\rangle$. The second condition for geometric convergence as given in Fact 2.4 tells us that there exist sequences $n_{i},m_{i}\in\mathbb{Z}$ such that $\psi_{i}^{n_{i}}\rightarrow a$ and $\psi_{i}^{m_{i}}\rightarrow b$ in $\text{PSL}_{2}\mathbb{C}$, as $i\rightarrow\infty$. Following the framework used above to create the injective homomorphisms $J_{s,t}$ in the setting of paths, one might hope to define an injective homomorphism $J_{i}:\langle a,b\rangle\rightarrow\text{PSL}_{2}\mathbb{C}$ for $i$ sufficiently large by setting $J_{i}(a)=\psi_{i}^{n_{i}}$ and $J_{i}(b)=\psi_{i}^{m_{i}}$. We can see, however, that such a map $J_{i}$ will never be an injective homomorphism.
Lemma 4.8.
For any $s\in[0,1]$ and subgroup $H\leq\Gamma_{s}$, $J_{s,t}(H)\cap\text{PSL}_{2}\mathbb{C}$ is a subgroup of $\Gamma_{t}$. Further, if $H$ is finitely generated, the map $\Psi:[0,1]\rightarrow\mathcal{D}$ given by $\Psi(t)=J_{s,t}(H)\cap\text{PSL}_{2}\mathbb{C}$ is a path in the geometric topology on $\mathcal{D}$. In particular, if $J_{s,t}|_{H}$ is injective for all $t\in I$, the map $J:I\rightarrow S(H)$ given by $J(t)=J_{s,t}|_{H}$ is a path in the topology of strong convergence.
Proof.
Let
$H^{\prime}=\{\psi\in H\;|\;t\in I_{\psi}^{s}\}.$
For all $\psi,\varphi\in H$, we have $I_{\psi}^{s}=I_{\psi^{-1}}^{s}$ and $I_{\psi}^{s}\cap I_{\varphi}^{s}\subseteq I_{\psi\varphi}^{s}$, so $H^{\prime}$ is a subgroup of $\Gamma_{s}$. Observe that $J_{s,t}(H)\cap\text{PSL}_{2}\mathbb{C}=J_{s,t}(H^{\prime})$ and Proposition 4.6 ($3\Rightarrow 2$) tells us that $J_{s,t}|_{H^{\prime}}$ is an injective homomorphism. The image of $J_{s,t}|_{H^{\prime}}$ lies in $\Gamma_{t}$, so we see $J_{s,t}(H^{\prime})\leq\Gamma_{t}$.
Now, assume that $H$ is finitely generated. We will show that $\Psi$ is continuous at a fixed $t\in[0,1]$. By combining Lemma 4.5(2) and Corollary 4.7, it will suffice to assume that $t\neq s$ and to consider a one sided limit. That is, we will assume $s<t$ (the other direction being handled analogously) and prove that
$J_{s,t^{\prime}}(H)\cap\text{PSL}_{2}\mathbb{C}\rightarrow J_{s,t}(H)\cap\text{PSL}_{2}\mathbb{C}$
as $t^{\prime}\rightarrow t$ from below.
We will check the two conditions for convergence in the geometric topology on $\mathcal{D}$, as stated in Fact 2.4. To check the first condition, suppose there exist sequences $\{t_{k}\}_{k\geq 1}\subseteq[s,t)$ and $\{\psi_{k}\}_{k\geq 1}\subseteq H$ such that $t_{k}\rightarrow t$ and $J_{s,t_{k}}(\psi_{k})\rightarrow\varphi\in\text{PSL}_{2}\mathbb{C}$ as $k\rightarrow\infty$. Since $\Gamma_{t_{k}}\rightarrow\Gamma_{t}$ as $k\rightarrow\infty$, we must have $\varphi\in\Gamma_{t}$. Letting $U_{\varphi}\subseteq\text{PSL}_{2}\mathbb{C}$ be the neighborhood of $\varphi$ guaranteed by Lemma 4.1 with respect to $\Gamma_{t}$, recall that $J_{t,t_{k}}(\varphi)$ is the unique element of $\Gamma_{t_{k}}$ contained in $U_{\varphi}$, for $k$ sufficiently large. Additionally, we see $J_{s,t_{k}}(\psi_{k})\in U_{\varphi}$ for $k$ sufficiently large, so potentially removing finitely many terms of the sequence and reindexing, we must have $J_{s,t_{k}}(\psi_{k})=J_{t,t_{k}}(\varphi)$ for all $k$. Note that it then follows that $[s,t]\subseteq I_{\psi_{k}}^{s}$ and $[t_{k},t]\subseteq I_{\varphi}^{t}$ for all $k$, so applying Lemma 4.5(2) twice tells us that
$J_{s,t}(\psi_{k})=(J_{t_{k},t}\circ J_{s,t_{k}})(\psi_{k})=J_{t_{k},t}(J_{t,t_{k}}(\varphi))=J_{t,t}(\varphi)=\varphi$
for all $k$. This confirms that $\varphi\in J_{s,t}(H)\cap\text{PSL}_{2}\mathbb{C}$, as was to be shown.
The second condition for geometric convergence follows from the definition of $J_{s,t}$, since every element of $J_{s,t}(H)\cap\text{PSL}_{2}\mathbb{C}$ is the limit of a path $J_{s,t^{\prime}}(\psi)$ for some $\psi\in H$, as $t^{\prime}\rightarrow t$. ∎
With a little more work, the hypothesis in the second statement of Lemma 4.8 that $H$ is finitely generated may be omitted, though we will not need this fact, so do not include the details.
4.3 Paths of Convex Cocompact manifolds
In Example C.2 above, if $G$ and $H$ are infinite covolume convex cocompact Kleinian groups, we obtain a path $\Gamma_{t}$ in $\mathcal{D}$ of convex cocompact Kleinian groups that split as a free product isomorphic to $G*H$, where one of the factors diverges to infinity in the limit. The goal for this section will be to understand generic paths of convex cocompact Kleinian groups. In particular, we will show that the phenomenon exhibited in Example C.2 is reflective of general paths of convex cocompact Kleinian groups: the only way the isomorphism type can change along such a path is for free factors to come in from or diverge to infinity. In particular, we will prove the following theorem.
Theorem C.
Suppose $\Gamma:[0,1]\rightarrow\mathcal{D}$ is a path such that $\Gamma(t)=\Gamma_{t}$ is convex cocompact for all $t$. Then, $H_{s}^{J}=J_{1,s}(\Gamma_{1})\cap\text{PSL}_{2}\mathbb{C}$ is a free factor of $\Gamma_{s}$ for all $s$. In particular, there exist finitely many subgroups $H_{1},...,H_{n}\leq\Gamma_{s}$ and times $s<t_{1}<...<t_{n}\leq 1$ so that
  1. 1.
    $\Gamma_{s}=H_{s}^{J}*H_{1}*...*H_{n}$ and
  2. 2.
    $J_{s,t}(\psi)\rightarrow\infty$ as $t\rightarrow t_{i}$, for all $\psi\in H_{i}$ and $1\leq i\leq n$.
Here, the maps $J_{s,t}:\Gamma_{s}\rightarrow\overline{\text{PSL}_{2}\mathbb{C}}$ for $s,t\in[0,1]$ are as defined in Equation 1. Since $\Gamma_{1}$ is finitely generated, Lemma 4.7 implies that $J_{1,t}$ is an injective homomorphism for $t$ close to $1$. For general $t$, however, $J_{1,t}$ need not be injective, nor a homomorphism. By Proposition 4.6, $J_{1,t}$ failing to be injective exactly corresponds to having $\infty\in J_{1,t}(\Gamma_{1})$. This case is why we consider $J_{1,s}(\Gamma_{1})\cap\text{PSL}_{2}\mathbb{C}$ in the statement, as this is always a subgroup of $\Gamma_{s}$, by Lemma 4.8.
We also note that every non-trivial splitting of the fundamental group of a convex cocompact hyperbolic $3$-manifold as a free product is induced by a separating properly embedded disk. As is discussed in [BEN, Section 2.4.1], this implies that any convex cocompact Kleinian group which splits as a nontrivial free product $\Gamma=H_{1}*H_{2}$ may be constructed from $H_{1}$ and $H_{2}$ by the first Klein-Maskit combination theorem. Thus, Theorem C completely generalizes the construction in Example C.2 and characterizes paths of convex cocompact Kleinian groups.
Example B.2 provides a sharpness to this result, showing that the conclusion does not hold if we replace ‘convex cocompact’ with ‘geometrically finite’ in the hypothesis of Theorem C; if $\Gamma_{1}$ contains a rank-$1$ cusp subgroup $\langle\psi\rangle$, Example B.2 constructs a path $t\mapsto\langle\Gamma_{1},\varphi_{t}\rangle$ converging to $\Gamma_{1}$ where $\psi$ commutes with $\varphi_{t}\not\in\Gamma_{1}$, and hence $\Gamma_{1}$ is not a free factor of $\langle\Gamma_{1},\varphi_{t}\rangle$ for any $t$. We expect, however, that with additional hypotheses, statements similar to Theorem C can be made in more generality.
Sequences of convex cocompact Kleinian groups converging to a convex cocompact limit are much less controlled than the analogous situation for paths of convex cocompact Kleinian group described by Theorem C. For example, any Kleinian group $\Gamma$ (in particular, any convex cocompact Kleinian group) may be approximated by a sequence of convex cocompact Kleinian groups $\{\Gamma_{n}\}$ which do not split as a non-trivial free product. To see this, apply Proposition 3.8 to choose a sequence $\Gamma_{n}^{\prime}\rightarrow\Gamma$ such that each $\partial CC(\mathbb{H}^{3}/\Gamma_{n}^{\prime})$ is a disjoint union of totally geodesic thrice punctured spheres. Each $\mathbb{H}^{3}/\Gamma_{n}^{\prime}$ is homeomorphic to the interior of a compact $3$-manifold with incompressible boundary. Example B.2 allows one to approximate each $\Gamma_{n}^{\prime}$ by geometrically finite Kleinian groups so that the corresponding hyperbolic $3$-manifolds have incompressible boundary and no rank-$1$ cusps. By performing Dehn surgery on any rank-$2$ cusps, these geometrically finite Kleinian groups may be further approximated by a sequence of convex cocompact Kleinian groups. Taking a diagonalization yields a sequence of convex cocompact Kleinian groups $\{\Gamma_{n}\}$ such that each $\mathbb{H}^{3}/\Gamma_{n}$ has incompressible boundary and $\Gamma_{n}\rightarrow\Gamma$.
Our first step towards proving Theorem C will be to develop a way to geometrically recognize a free factor of the fundamental group of a hyperbolic $3$-manifold. To do this, we will need some basic facts about loxodromic isometries in $\text{PSL}_{2}\mathbb{C}$. Each loxodromic isometry $\psi\in\text{PSL}_{2}\mathbb{C}$ preserves a unique geodesic $g_{\psi}$ in $\mathbb{H}^{3}$, its axis of translation, on which $\psi$ acts by translation by some length $\ell(\psi)$. Note that $\ell(\psi)=\min_{x\in\mathbb{H}^{3}}d_{\mathbb{H}^{3}}(x,\psi\cdot x)$ and $\ell(\psi)$ is a continuous function of the trace $\text{tr}(\psi)$. For a loxodromic isometry $\psi$ we define
$T(\psi)=\{x\in\mathbb{H}^{3}\;|\;d_{\mathbb{H}^{3}}(x,\psi\cdot x)\leq 1\}.$
Additionally, for $M$ a hyperbolic $3$-manifold, a subset $A\subseteq M$, and $D>0$, let
$\mathcal{N}_{D}(A)=\{x\in M:d_{M}(x,A)\leq D\}$
be the closed radius $D$ neighborhood of $A$.
Lemma 4.9.
Let $\psi\in\text{PSL}_{2}\mathbb{C}$ be a loxodromic isometry.
  1. 1.
    There exists $R=R(\ell(\psi))$ such that $T(\psi)\subseteq\mathcal{N}_{R}(g_{\psi})$, and $R(\ell(\psi))$ may be taken to be a decreasing function of $\ell(\psi)$.
  2. 2.
    For all $\eta>0$, there exists $D=D(\eta)$ such that for all $x\in\mathbb{H}^{3}$ satisfying $d_{\mathbb{H}^{3}}(x,\psi\cdot x)\leq\eta$, we have $d_{\mathbb{H}^{3}}(x,T(\psi)\cup g_{\psi})\leq D$.
Importantly, the value $D$ in Lemma 4.9(2) only depends on $\eta$, and does not depend on $\psi$. Additionally, in Lemma 4.9(2) we consider the distance from $x$ to the set $T(\psi)\cup g_{\psi}$, since $T(\psi)=\varnothing$ if $\ell(\psi)>1$.
Proof of Lemma 4.9:.
(1): If $\ell(\psi)>1$, take $R=0$. Suppose, then, that $\ell(\psi)\leq 1$. Note that $T(\psi)$ is a metric neighborhood of $g_{\psi}$, the axis of translation of $\psi$; that is, there exists $R_{\psi}$ so that $T(\psi)=\mathcal{N}_{R_{\psi}}(g_{\psi})$. Further, $R_{\psi}$ will be maximized in the case where $\psi$ has no rotation part (that is, $\psi$ has real trace). More precisely, fix a loxodromic $\varphi\in\text{PSL}_{2}\mathbb{C}$ with $\ell(\varphi)=\ell(\psi)$ such that $\varphi$ has no rotation part; then $R_{\psi}\leq R_{\varphi}$. Noting that $R_{\varphi}$ depends only on $\ell(\varphi)=\ell(\psi)$, this immediately gives existence of the value $R=R(\ell(\psi))$ by setting $R=R_{\varphi}$. It will be useful for the second statement in the lemma to explicitly compute $R_{\varphi}$, so we do this now.
Figure 4: This depicts the Saccheri quadrilateral constructed in the proof of Lemma 4.9.
Fix $x\in\partial T(\varphi)$, so that $d_{\mathbb{H}^{3}}(x,\varphi\cdot x)=1$, and note that
$d_{\mathbb{H}^{3}}(x,g_{\varphi})=d_{\mathbb{H}^{3}}(\varphi\cdot x,g_{\varphi})=R_{\varphi}.$
Since $\varphi$ has no rotation part, there exists a $2$-dimensional geodesic hyperplane containing $g_{\varphi}$, $x$, and $\varphi\cdot x$. Let $\overline{x}$ and $\overline{\varphi\cdot x}$ denote the orthogonal projections to $g_{\varphi}$ of $x$ and $\varphi\cdot x$, respectively. Observe that the points $x$, $\varphi\cdot x$, $\overline{\varphi\cdot x}$, $\overline{x}$ form a Saccheri quadrilateral, as shown in Figure 4. Since $d_{\mathbb{H}^{3}}(\overline{x},\overline{\varphi\cdot x})=d_{\mathbb{H}^{3}}(\overline{x},\varphi\cdot\overline{x})=\ell(\varphi)$, a standard computation gives that
$R_{\varphi}=\operatorname{arcosh}\left(\frac{\sinh(1/2)}{\sinh(\ell(\varphi)/2)}\right).$ (2)
(2): Define
$T_{\eta}(\psi)=\{x\in\mathbb{H}^{3}\;|\;d_{\mathbb{H}^{3}}(x,\psi\cdot x)\leq\eta\}.$
If $\eta<1$, then $T_{\eta}(\psi)\subseteq T(\psi)$, so we may set $D=0$ in this case. Assume that $\eta\geq 1$. Then, it suffices to consider $\psi$ such that $\ell(\psi)\leq\eta$, else $T_{\eta}(\psi)=\varnothing$. As in the proof of the first statement above, recall that in the case $\ell(\psi)\leq 1$, there exists $R_{\psi}\geq 0$ so that $T(\psi)=\mathcal{N}_{R_{\psi}}(g_{\psi})$. Setting $R_{\psi}=0$ if $\ell(\psi)>1$, it follows that there exists $D_{\psi}\geq 0$ so that
$T_{\eta}(\psi)=\mathcal{N}_{R_{\psi}+D_{\psi}}(g_{\psi})=\mathcal{N}_{D_{\psi}}(T_{\psi}\cup g_{\psi}).$
Fix a loxodromic $\varphi\in\text{PSL}_{2}\mathbb{C}$ with $\ell(\varphi)=\ell(\psi)$ so that $\varphi$ has no rotation part, and note that $D_{\psi}\leq D_{\varphi}$. Our goal will be to show that there exists an upper bound for $D_{\varphi}$ that depends only on $\eta$.
Note that $D_{\varphi}$ is determined by $\ell(\varphi)$. For $1\leq\ell(\varphi)\leq\eta$, observe that $D_{\varphi}$ is a decreasing function of $\ell(\varphi)$. Therefore, it suffices to assume $\ell(\varphi)\leq 1$ and show that $D_{\varphi}$ is uniformly bounded in this case. To compute $D_{\varphi}$, choose $y\in\partial T_{\eta}(\varphi)$, so that $d_{\mathbb{H}^{3}}(y,\varphi\cdot y)=\eta$ and
$d_{\mathbb{H}^{3}}(y,g_{\varphi})=d_{\mathbb{H}^{3}}(\varphi\cdot y,g_{\varphi})=R_{\varphi}+D_{\varphi}.$
Assume that the point $x$ from the proof of the first statement lies on the geodesic segment orthogonal to $g_{\varphi}$ through $y$. Then, the points $y$, $\varphi\cdot y$, $\overline{\varphi\cdot x}$, $\overline{x}$ form a Saccheri quadrilateral, as shown in Figure 4. Thus, a standard computation shows that
$D_{\varphi}=\operatorname{arcosh}\left(\frac{\sinh(\eta/2)}{\sinh{(\ell(\varphi)/2)}}\right)-R_{\varphi}=\operatorname{arcosh}\left(\frac{\sinh(\eta/2)}{\sinh{(\ell(\varphi)/2)}}\right)-\operatorname{arcosh}\left(\frac{\sinh(1/2)}{\sinh(\ell(\varphi)/2)}\right)$
where the second equality is given by Equation 2. To show that $D_{\varphi}$ is bounded, it will suffice to show that $\lim_{\ell(\varphi)\rightarrow 0}D_{\varphi}$ exists. Note that for any functions $f,g$ such that $\lim_{x\rightarrow 0}f(x)=\lim_{x\rightarrow 0}g(x)=\infty$ and $\lim_{x\rightarrow 0}(\ln f(x)-\ln g(x))$ exists, one has
$\lim_{x\rightarrow 0}(\operatorname{arcosh}f(x)-\operatorname{arcosh}g(x))=\lim_{x\rightarrow 0}(\ln f(x)-\ln g(x)).$
Thus, since
$\lim_{\ell(\varphi)\rightarrow 0}\ln\left(\frac{\sinh(\eta/2)}{\sinh{(\ell(\varphi)/2)}}\right)-\ln\left(\frac{\sinh(1/2)}{\sinh(\ell(\varphi)/2)}\right)=\ln(\eta/2)-\ln(1/2)$
we confirm that $D_{\varphi}$ is uniformly bounded in terms of $\eta$, and the claim follows. ∎
For a Kleinian group $\Gamma$, we define
$T(\Gamma)=\bigcup_{\psi\in\Gamma}T(\psi)=\{x\in\mathbb{H}^{3}\;|\;\exists\psi\in\Gamma\text{ such that }d_{\mathbb{H}^{3}}(x,\psi\cdot x)\leq 1\}.$
Observe that $T(\Gamma)$ is the $\Gamma$-lift of the $1$-thin part of $\mathbb{H}^{3}/\Gamma$. We then make the following definition, which is a weak version of one made in [IANRANK].
Definition 4.10.
For a Kleinian group $\Gamma$, the (weak) thick convex hull of $\Gamma$ is the set $TCH(\Gamma)=CH(\Gamma)\cup T(\Gamma)$. The (weak) thick convex core of $M=\mathbb{H}^{3}/\Gamma$, denoted $TCC(M)$, is the projection of $TCH(\Gamma)$ to $M$.
Note that $TCC(M)=CC(M)\cup M_{\leq 1}$. The next lemma gives the key property which we will need of the thickened convex core. Roughly, since we have included all thin parts in $TCC(M)$, all points in $M$ of bounded injectivity radius are within a uniformly bounded distance of $TCC(M)$.
Lemma 4.11.
For all $\eta>0$, there exists $D=D(\eta)>0$ so that if $M$ is a convex cocompact hyperbolic $3$-manifold, $M_{\leq\eta}\subseteq\mathcal{N}_{D}(TCC(M))$.
Proof.
Let $D=D(\eta)$ be the constant given by Lemma 4.9(2) with respect to $\eta$. Suppose $x\in\mathbb{H}^{3}$ satisfies $d_{\mathbb{H}^{3}}(x,\psi\cdot x)\leq\eta$ for some $\psi\in\Gamma$. Noting that $T(\psi)\cup g_{\psi}\subseteq TCH(\Gamma)$, it follows that $d_{\mathbb{H}^{3}}(x,TCH(\Gamma))\leq D$. ∎
Fix $M$ a convex cocompact hyperbolic $3$-manifold, and fix $\xi>0$ less than the $2$ and $3$-dimensional Margulis constants. Consider a component, $A$, of the $\xi$-thin part of the surface $\partial CC(M)$, $(\partial CC(M))_{\leq\xi}$. $A$ is an annulus with boundary curves $\alpha_{1}$ and $\alpha_{2}$ of length between $\xi$ and $2\xi$. Suppose that $\alpha_{1}$ (and hence $\alpha_{2}$) is trivial in $M$. Then, each $\alpha_{i}$ bounds a disk $D_{i}\subseteq CC(M)$ of diameter at most $\xi$ with $D_{1}\cap D_{2}=\varnothing$. The $2$-sphere $D_{1}\cup A\cup D_{2}$ in $M$ must bound a ball $h\subseteq CC(M)$. After Bowditch in [BOW], we will call $h$ a $\xi$-handle. Note that we can assume that every point in $h$ is contained in a compressing disk $(D,\partial D)\subseteq(h,A)$ with diameter at most $\xi$. By the choice of $\xi$, all of the $\xi$-handles constructed as above can be made disjoint. We let $W=W(\xi)\subseteq CC(M)$ denote the union of all $\xi$-handles.
The main result of Bowditch in [BOW] is that there exists $\eta>0$ depending only on $\xi$ and the topology of $M$ such that $CC(M)\subseteq M_{\leq\eta}\cup W$. The following proposition provides a way of recognizing a free factor of the fundamental group of a hyperbolic $3$-manifold.
Proposition 4.12.
Let $\Gamma\in\mathcal{D}$ be a convex cocompact Kleinian group, and fix $H\leq\Gamma$. For $M=\mathbb{H}^{3}/\Gamma$ and $M_{H}=\mathbb{H}^{3}/H$, let $P:M_{H}\rightarrow M$ denote the covering map. Then, there exists $D>0$ depending only on the topology of $M$ so that if $\mathcal{N}_{D}(TCC(M_{H}))$ isometrically embeds into $M$ under $P$, then $H$ is a free factor of $\Gamma$.
Proof.
Let $\eta>0$ be the constant given by Bowditch in [BOW] so that $CC(M)\subseteq M_{\leq\eta}\cup W$. As in Lemma 4.11, fix $D^{\prime}>0$ so that $(M_{H})_{\leq\eta+1}\subseteq\mathcal{N_{D^{\prime}}}(TCC(M_{H}))$. Set $D=D^{\prime}+4\xi+\eta+1$ and suppose that $P$ embeds $\mathcal{N}_{D}(TCC(M_{H}))$ isometrically into $M$.
Set
$N=\{p\in M_{H}\;|\;D^{\prime}\leq d_{M_{H}}(p,TCC(M_{H}))\leq D^{\prime}+4\xi\}$
and for $0\leq j\leq 4$ let
$S_{j}=\{p\in M_{H}\;|\;d_{M_{H}}(p,TCC(M_{H}))=D^{\prime}+j\cdot\xi\}.$
Note that $N$ separates $M_{H}$ and $P(N)$ separates $M$. By the choice of $D^{\prime}$, $\text{inj}_{M_{H}}(q)>\eta$ for all $q\in N$; for such $q$, since $P$ isometrically embeds $B_{M_{H}}(q,\eta+1)$ into $M$, the injectivity radius in $M$ is more than $\eta$ at all points of $P(N)$. Thus, $CC(M)$ may only intersect $P(N)$ in $W$, by the choice of $\eta$.
Let $h_{0}$ be a component of a $P(N)\cap W$ such that $h_{0}$ intersects both $P(S_{0})$ and $P(S_{4})$. Since $P(N)$ separates $M$, $h_{0}$ must also intersect $P(S_{2})$. Choose $x\in h_{0}\cap P(S_{2})$ and recall that we may fix a compressing disk $(D_{0},\partial D_{0})\subseteq(h_{0},\partial W)$ of diameter at most $\xi$ which contains $x$.
Let $C\subseteq W$ be the (disjoint) union of all such compressing disks $D_{0}$, each constructed as above from a unique component $h_{0}$ of $P(N)\cap W$ which intersects both $P(S_{0})$ and $P(S_{4})$. Let $X$ be the component of $CC(M)\backslash C$ containing $P(CC(M_{H}))$. Fix $p^{\prime}\in CC(M_{H})$ and let $p=P(p^{\prime})$. Observe that $\pi_{1}(X,p)$ is a free factor of $\pi_{1}(CC(M),p)$, since $X$ is a component of the result of removing a disjoint union of compressing disks from $CC(M)$.
We now claim that
$X\subseteq P(\mathcal{N}_{D}(TCC(M_{H}))).$ (3)
To see this, suppose $\gamma:[0,1]\rightarrow CC(M)$ is a path with $\gamma(0)\in P(CC(M_{H}))$ and $\gamma(1)\not\in P(\mathcal{N}_{D}(TCC(M_{H})))$. Since $P(N)$ separates $M$, there must be $t_{1},t_{2}\in[0,1]$ such that
$\gamma(t_{1})\in P(S_{0}),\;\;\gamma(t_{2})\in P(S_{4}),\;\text{ and }\;\gamma(t)\in P(N)\text{ for all }t\in[t_{1},t_{2}].$
Then, there must be some component $h_{0}$ of $P(N)\cap W$, as above, so that $\gamma([t_{1},t_{2}])\subseteq h_{0}$. Hence, we see that the disk $D_{0}$ separates $\gamma(t_{1})$ and $\gamma(t_{2})$ in $h_{0}$, which confirms the claim in Equation 3.
Where $p^{\prime}\in CC(M_{H})$ is as above, any loop $\beta^{\prime}:[0,1]\rightarrow M_{H}$ with $\beta^{\prime}(0)=\beta^{\prime}(1)=p^{\prime}$ is pointed homotopic into $CC(M_{H})$. Thus, since $X\subseteq P(\mathcal{N}_{D}(TCC(M_{H})))$, any loop $\beta:[0,1]\rightarrow X$ with $\beta(0)=\beta(1)=p$ is pointed homotopic in $M$ into $P(CC(M_{H}))$. Therefore, we see that $\pi_{1}(X,p)=\pi_{1}(P(CC(M_{H})),p)$. Let $P_{*}:\pi_{1}(M_{H},p^{\prime})\rightarrow\pi_{1}(M,p)$ be the injective homomorphism induced from the projection map $P$, and note that
$P_{*}(\pi_{1}(M_{H},p^{\prime}))=\pi_{1}(P(CC(M_{H})),p)=\pi_{1}(X,p).$
Thus, we have shown that $P_{*}(\pi_{1}(M_{H},p^{\prime}))$ is a free factor of $\pi_{1}(CC(M),p)=\pi_{1}(M,p)$.
Fix an $H$-lift $\tilde{p}\in\mathbb{H}^{3}$ of $p^{\prime}\in CC(M_{H})$ and let $\Psi_{H}:\pi_{1}(M_{H},p^{\prime})\rightarrow H$ be the holonomy representation determined by $\tilde{p}$ and $H$. Note that $\tilde{p}$ is also a $\Gamma$-lift of $p\in M$, so we may let $\Psi:\pi_{1}(M,p)\rightarrow\Gamma$ be the holonomy representation determined by $\tilde{p}$ and $\Gamma$. Note that the following diagram commutes.
${{\pi_{1}(M_{H},p^{\prime})}}$${{\pi_{1}(M,p)}}$${H}$${\Gamma}$$\scriptstyle{P_{*}}$$\scriptstyle{\Psi_{H}}$$\scriptstyle{\Psi}$inclusion
Since the vertical maps in the diagram are isomorphisms, we have confirmed that $H$ is a free factor of $\Gamma$. ∎
The general strategy for proving Theorem C will be to use Proposition 4.12 to justify that $J_{1,t}(\Gamma_{t})$ is a free factor of $\Gamma_{t}$, for a path $t\mapsto\Gamma_{t}$ as in the theorem. To apply Proposition 4.12, we will first need to be able to control $TCH(J_{1,t}(\Gamma_{1}))$ for $t$ close to $1$. The next two lemmas will address this.
For the next lemma, it will be useful to have the language of Chabauty convergence of closed sets in $\mathbb{H}^{3}$. Letting $\mathcal{C}(\mathbb{H}^{3})$ denote the set of closed subsets of $\mathbb{H}^{3}$, the conditions for convergence in the Chabauty topology on $\mathcal{C}(\mathbb{H}^{3})$ are completely analogous to those for convergence in the geometric topology on $\mathcal{D}$, as described in Fact 2.4. In particular, we have $X_{n}\rightarrow X$ in $\mathcal{C}(\mathbb{H}^{3})$ if (i) every accumulation point of a sequence $x_{n}\in X_{n}$ is contained in $X$, and (ii) every $x\in X$ is the limit of a sequence $x_{n}\in X_{n}$. See [BEN, Chapter E] for more information on the Chabauty topology.
Additionally, for a convex cocompact Kleinian group $\Gamma$, we make the following definitions. Let $\mathcal{P}(\Gamma)$ denote the Dirichlet fundamental polyhedron for $\Gamma$ based at our fixed basepoint $O\in\mathbb{H}^{3}$ (see [BEN, C.1.3]). We also let $\text{sys}(\Gamma)$ denote the systole of $\mathbb{H}^{3}/\Gamma$; that is, $\text{sys}(\Gamma)$ is the length of the shortest geodesic in $\mathbb{H}^{3}/\Gamma$, which is given by
$\text{sys}(\Gamma)=\min_{\psi\in\Gamma}\ell(\psi).$
Lemma 4.13.
Let $\Gamma$ be a convex cocompact Kleinian group and fix $\rho\in GF(\Gamma)$. Then, for any sequence $\{\rho_{n}\}\subseteq GF(\Gamma)$ converging strongly to $\rho$,
  1. 1.
    there exists a compact set $K\subseteq\mathbb{H}^{3}$ so that $CH(\rho_{n}(\Gamma))\subseteq\rho_{n}(\Gamma)\cdot K$ for $n$ sufficiently large,
  2. 2.
    $\text{sys}(\rho_{n}(\Gamma))\rightarrow\text{sys}(\rho(\Gamma))$ as $n\rightarrow\infty$,
  3. 3.
    there exists $R>0$ such that $T(\rho_{n}(\Gamma))\subseteq\mathcal{N}_{R}(CH(\rho_{n}(\Gamma)))$ for $n$ sufficiently large,
  4. 4.
    for all $D\geq 0$, there exists a compact set $K^{\prime}\subseteq\mathbb{H}^{3}$ such that $\mathcal{N}_{D}(TCH(\rho_{n}(\Gamma)))\subseteq\rho_{n}(\Gamma)\cdot K^{\prime}$ for $n$ sufficiently large.
Additionally, $K$ and $R$ depend only on $\rho$, and $K^{\prime}$ depends only on $\rho$ and $D$.
Proof.
(1): Define
$A_{n}=\mathcal{P}(\rho_{n}(\Gamma))\cap CH(\rho_{n}(\Gamma)),\;\;\;\;\;\;A=\mathcal{P}(\rho(\Gamma))\cap CH(\rho(\Gamma)).$
Fixing $\varepsilon>0$ and letting $K=\mathcal{N}_{\varepsilon}(A)$, it will suffice to show that $A_{n}\subseteq K$ for $n$ sufficiently large.
A result of Jørgensen and Marden [JORMAR] implies that the limit sets $\Lambda(\rho_{n}(\Gamma))$ converge to $\Lambda(\rho(\Gamma))$ in the Hausdorff topology on $S^{2}_{\infty}$. This, combined with a result of Bowditch on convex hulls [BOWCHS], implies that $CH(\rho_{n}(\Gamma))\rightarrow CH(\rho(\Gamma))$ in the Chabauty topology on $\mathcal{C}(\mathbb{H}^{3})$. Jørgensen and Marden [JORMAR] also establish that strong convergence implies convergence of the Dirichlet fundamental polyhedra. More precisely, it follows that for all $R>0$
$\overline{B_{\mathbb{H}^{3}}(O,R)}\cap\mathcal{P}(\rho_{n}(\Gamma))\rightarrow\overline{B_{\mathbb{H}^{3}}(O,R)}\cap\mathcal{P}(\rho(\Gamma))$
in the Hausdorff topology on compact sets of $\mathbb{H}^{3}$. This implies that $\mathcal{P}(\rho_{n}(\Gamma))\rightarrow\mathcal{P}(\rho(\Gamma))$ in the Chabauty topology on $\mathcal{C}(\mathbb{H}^{3})$.
First, we will show there exists a sequence of points $p_{n}\in A_{n}$ converging in $\mathbb{H}^{3}$ to a point $p\in A$. Choose $p\in A$ so that $p$ is on the interior of $\mathcal{P}(\rho(\Gamma))$. By the Chabauty convergence $CH(\rho_{n}(\Gamma))\rightarrow CH(\rho(\Gamma))$, there exists a sequence $p_{n}\in CH(\rho_{n}(\Gamma))$ with $p_{n}\rightarrow p$. We see from the convergence of Dirichlet polyhedra discussed in the preceding paragraph that $p_{n}$ must be contained in the interior of $\mathcal{P}(\rho_{n}(\Gamma))$ for $n$ sufficiently large, so $p_{n}\in A_{n}$ for $n$ sufficiently large.
Now, suppose for contradiction that there exists a sequence $x_{n}\in A_{n}$ with $d_{\mathbb{H}^{3}}(x_{n},A)>\varepsilon$ for all $n$. By potentially replacing each $x_{n}$ with a point on the geodesic segment between $p_{n}$ and $x_{n}$, the convexity of $A_{n}$ and the compactness of $A$ allows us to assume that the sequence $\{x_{n}\}$ is bounded. Hence, the sequence $\{x_{n}\}$ has an accumulation point $x\in\mathbb{H}^{3}$, so the Chabauty convergence of the sequences $CH(\rho_{n}(\Gamma))\rightarrow CH(\rho(\Gamma))$ and $\mathcal{P}(\rho_{n}(\Gamma))\rightarrow\mathcal{P}(\rho(\Gamma))$ implies that $x\in A$. It must also be the case that $d_{\mathbb{H}^{3}}(x,A)\geq\varepsilon$, so we have a contradiction.
(2): First, pick $\psi\in\Gamma$ so that $\ell(\rho(\psi))=\text{sys}(\rho(\Gamma))$. Since $\rho_{n}(\psi)\rightarrow\rho(\psi)$ in $\text{PSL}_{2}\mathbb{C}$, we have $\ell(\rho_{n}(\psi))\rightarrow\ell(\rho(\psi))$ as $n\rightarrow\infty$. It then follows that
$\limsup_{n}\text{sys}(\rho_{n}(\Gamma))\leq\text{sys}(\rho(\Gamma)).$ (4)
By the proof of statement (1) of the lemma, there exists a compact set $K\subseteq\mathbb{H}^{3}$ so that $K$ contains a fundamental domain for the action of $\rho_{n}(\Gamma)$ on $CH(\rho_{n}(\Gamma))$ for sufficiently large $n$. For such $n$, there is some $\psi_{n}\in\rho_{n}(\Gamma)$ so that $\ell(\psi_{n})=\text{sys}(\rho_{n}(\Gamma))$ and $g_{\psi_{n}}$ intersects $K$. Let $K^{\prime\prime}=\mathcal{N}_{\text{sys}(\rho(\Gamma))}(K)$. The preceding paragraph tells us that $\text{sys}(\rho_{n}(\Gamma))<2\text{sys}(\rho(\Gamma))$ for $n$ sufficiently large; hence, for such $n$, $K^{\prime\prime}$ contains a geodesic segment $\alpha_{n}\subseteq g_{\psi_{n}}$ so that $g_{\psi_{n}}=\langle\psi_{n}\rangle\cdot\alpha_{n}$. In particular, we see $\psi_{n}\cdot K^{\prime\prime}\cap K^{\prime\prime}\neq\varnothing$ for $n$ sufficiently large so it follows that the sequence $\{\psi_{n}\}_{n}\subseteq\text{PSL}_{2}\mathbb{C}$ is bounded.
Now, choose a subsequence $\{n_{i}\}_{i}$ so that
$\liminf_{n}\text{sys}(\rho_{n}(\Gamma))=\lim_{i\rightarrow\infty}\text{sys}(\rho_{n_{i}}(\Gamma))=\lim_{i\rightarrow\infty}\ell(\psi_{n_{i}}).$
Since the sequence $\{\psi_{n_{i}}\}_{i}\subseteq\text{PSL}_{2}\mathbb{C}$ is bounded, there is a further subsequence, which we denote the same way, converging to some $\psi\in\text{PSL}_{2}\mathbb{C}$. Since $\rho_{n}(\Gamma)\rightarrow\rho(\Gamma)$ in $\mathcal{D}$, the first condition for geometric convergence in $\mathcal{D}$ in Fact 2.4 implies that we must have $\psi\in\rho(\Gamma)$. We then conclude that
$\text{sys}(\rho(\Gamma))\leq\ell(\psi)=\lim_{i\rightarrow\infty}\ell(\psi_{n_{i}})=\liminf_{n}\text{sys}(\rho_{n}(\Gamma)).$ (5)
Combining Equations 4 and 5 gives the claim.
(3): Fix $\delta$ with $0<\delta<\text{sys}(\rho(\Gamma))$. Then, statement (2) of the lemma implies that $\text{sys}(\rho_{n}(\Gamma))\geq\delta$ for $n$ sufficiently large. Let $R=R(\delta)$ be the constant given by Lemma 4.9(1) with respect to translation length $\delta$. For large $n$ and any $\psi\in\rho_{n}(\Gamma)$, we then have $\ell(\psi)\geq\delta$, so $T(\psi)\subseteq\mathcal{N}_{R}(g_{\psi})$. Since $g_{\psi}\subseteq CH(\rho_{n}(\Gamma))$, we have confirmed that $T(\rho_{n}(\Gamma))\subseteq\mathcal{N}_{R}(CH(\rho_{n}(\Gamma)))$.
(4): Define $K^{\prime}=\mathcal{N}_{R+D}(K)$, where $K\subseteq\mathbb{H}^{3}$ and $R>0$ are as in statements (1) and (3) of the lemma, respectively. Then, we see
$\mathcal{N}_{D}(TCH(\rho_{n}(\Gamma)))\subseteq\mathcal{N}_{R+D}(CH(\rho_{n}(\Gamma)))\subseteq\rho_{n}(\Gamma)\cdot\mathcal{N}_{R+D}(K)$
for $n$ sufficiently large, so the result follows. ∎
Lemma 4.14.
Suppose $\Gamma:[0,1]\rightarrow\mathcal{D}$ is a path in the geometric topology. Setting $\Gamma(t)=\Gamma_{t}$, suppose that $\Gamma_{1}$ is convex cocompact. Set $M_{t}=\mathbb{H}^{3}/\Gamma_{t}$, $M_{t}^{J}=\mathbb{H}^{3}/J_{1,t}(\Gamma_{1})$, and let $P_{t}:M_{t}^{J}\rightarrow M_{t}$ be the induced covering map. Then, for any $D>0$, $\mathcal{N}_{D}(TCC(M_{t}^{J}))$ isometrically embeds into $M_{t}$ under $P_{t}$, for $t$ sufficiently close to $1$.
Proof.
By Corollary 4.7, $J_{1,t}$ is an injective homomorphism into $\text{PSL}_{2}\mathbb{C}$ for $t$ sufficiently close to $1$. Additionally, Marden’s stability theorem [MAR] then implies that $J_{1,t}(\Gamma_{1})$ is convex cocompact for $t$ sufficiently close to $1$, and hence for such $t$ we have $J_{1,t}\in GF(\Gamma_{1})$. Thus, taking the compact set $K^{\prime}\subseteq\mathbb{H}^{3}$ given by Lemma 4.13, there exists $t_{0}\in[0,1)$ so that $\mathcal{N}_{D}(TCH(J_{1,t}(\Gamma_{1})))\subseteq J_{1,t}(\Gamma_{1})\cdot K^{\prime}$ for all $t\geq t_{0}$.
Now, suppose there exists a sequence $\{t_{n}\}\subseteq[t_{0},1]$ with $t_{n}\rightarrow 1$ and elements $\psi_{n}\in\Gamma_{t_{n}}$ so that $\psi_{n}\cdot K^{\prime}\cap K^{\prime}\neq\varnothing$ for all $n$. Observe that the sequence $\{\psi_{n}\}\subseteq\text{PSL}_{2}\mathbb{C}$ must be bounded. Hence, passing to a subsequence, there exists $\psi\in\text{PSL}_{2}\mathbb{C}$ so that $\psi_{n}\rightarrow\psi$. Since $\Gamma_{t}\rightarrow\Gamma_{1}$ in $\mathcal{D}$, the first condition for geometric convergence of Kleinian groups in Fact 2.4 tells us that $\psi\in\Gamma_{1}$.
Lemma 4.1 implies that there exists a neighborhood $U_{\psi}\subseteq\text{PSL}_{2}\mathbb{C}$ so that for some $t_{0}^{\prime}\geq t_{0}$, $|\Gamma_{t}\cap U_{\psi}|=1$ for all $t\geq t_{0}^{\prime}$. Recall that for $t\geq t_{0}^{\prime}$, $J_{1,t}(\psi)$ is defined to be the unique element in $\Gamma_{t}\cap U_{\psi}$. Since $\psi_{n}\rightarrow\psi$, we must have $\psi_{n}\in U_{\psi}$ and therefore $\psi_{n}=J_{1,t_{n}}(\psi)$, for $n$ sufficiently large.
Thus, we conclude that for $t$ sufficiently close to $1$, if $\varphi\in\Gamma_{t}$ satisfies $\varphi\cdot K^{\prime}\cap K^{\prime}\neq\varnothing$, then we must have $\varphi\in J_{1,t}(\Gamma_{1})$. Hence, it follows that for such $t$ and all $\varphi^{\prime}\in\Gamma_{t}\backslash J_{1,t}(\Gamma_{1})$, we must have
$\varphi^{\prime}\cdot\mathcal{N}_{D}(TCH(J_{1,t}(\Gamma_{1})))\cap\mathcal{N}_{D}(TCH(J_{1,t}(\Gamma_{1})))=\varnothing$
which implies the claim. ∎
Lemma 4.15.
Suppose $\Gamma:[0,1]\rightarrow\mathcal{D}$ is a path in the geometric topology such that $\Gamma_{t}=\Gamma(t)$ is convex cocompact for all $t$ with $J_{0,t}(\Gamma_{0})=\Gamma_{t}$ for all $t<1$. Then, $J_{1,t}(\Gamma_{1})$ is a free factor of $\Gamma_{t}$ for all $t$. In particular, if $J_{0,1}$ is not finite valued, there exist a nontrivial subgroup $H_{0}\leq\Gamma_{0}$ so that $\Gamma_{t}$ splits as a free product $\Gamma_{t}=J_{1,t}(\Gamma_{1})*J_{0,t}(H_{0})$ for all $t<1$, with $J_{1,t}(\Gamma_{1})\cong\Gamma_{1}$ and $J_{0,t}(H_{0})\cong H_{0}$.
Recall from Section 4.2 that $J_{0,1}:\Gamma_{0}\rightarrow\overline{\text{PSL}_{2}\mathbb{C}}$ is said to be finite valued if its image is contained in $\text{PSL}_{2}\mathbb{C}$. In the case where $J_{0,1}$ is not finite valued, the subgroup $J_{0,t}(H_{0})$ and must diverge to infinity as $t\rightarrow 1$, by Proposition 4.6.
Proof of Lemma 4.15:.
Note that the hypothesis that $J_{0,t}(\Gamma_{0})=\Gamma_{t}$ for all $t<1$ implies that $J_{s,t}$ is an isomorphism between $\Gamma_{s}$ and $\Gamma_{t}$ for all $s,t<1$, by Proposition 4.6. Corollary 4.7 implies that $J_{1,t}$ is an injective homomorphism for $t$ close to $1$; hence, since $J_{s,t}$ is finite valued for all $s,t<1$, Proposition 4.6 implies that $I_{\psi}^{1}=[0,1]$ for all $\psi\in\Gamma_{1}$. Thus, Proposition 4.6 also implies that $J_{1,t}$ is an injective homomorphism for all $t$, and therefore that $J_{1,t}(\Gamma_{1})\cong\Gamma_{1}$ for all $t$.
Additionally, observe that $t\mapsto J_{0,t}$ is a path in $GF(\Gamma_{0})$ for $t\in[0,1)$ and that $M_{t}=\mathbb{H}^{3}/\Gamma_{t}$ is therefore homeomorphic to $M_{0}$ for all $t<1$. Let $D>0$ be the constant given by Proposition 4.12 with respect to the topology of $M_{0}$. Setting $M_{t}^{J}=\mathbb{H}^{3}/J_{1,t}(\Gamma_{1})$, there exists $t_{0}<1$ so that $\mathcal{N}_{D}(TCC(M_{t_{0}}^{J}))$ isometrically embeds into $M_{t_{0}}$ under the induced covering map $M_{t_{0}}^{J}\rightarrow M_{t_{0}}$, by Lemma 4.14. Thus, by Proposition 4.12, $J_{1,t_{0}}(\Gamma_{1})$ is a free factor of $\Gamma_{t_{0}}$. Since $J_{t_{0},0}$ is an isomorphism onto $\Gamma_{0}$, Lemma 4.5(2) implies that $J_{1,0}(\Gamma_{1})$ is a free factor of $\Gamma_{0}$. Thus, we may write $\Gamma_{0}=J_{1,0}(\Gamma_{1})*H_{0}$. Similarly, since each $J_{0,t}$ is an isomorphism onto $\Gamma_{t}$ for all $t<1$, $J_{0,t}(H_{0})\cong H_{0}$ and $\Gamma_{t}=J_{1,t}(\Gamma_{0})*J_{0,t}(H_{0})$ for all $t<1$.
Finally, note that if $J_{0,1}$ is not finite valued, then Proposition 4.6 says that $1\not\in I_{\psi}^{0}$ for some $\psi\in\Gamma_{0}$. In particular, $\psi\not\in J_{1,0}(\Gamma_{1})$, so $J_{1,0}(\Gamma_{1})$ must be a proper subgroup of $\Gamma_{0}$, implying that $H_{0}$ is a non-trivial subgroup. ∎
Finally, we can complete the proof of Theorem C.
Proof of Theorem C:.
Set $t_{0}=1$ and $G_{0}=\Gamma_{s}$. Inductively, assume that $t_{k}\neq 1$ and set
$t_{k+1}=\sup\{t\in[0,1]\;|\;\text{the restriction }J_{s,t}|_{G_{k}}\text{ is finite valued}\}.$
Imprecisely, we obtain $t_{k+1}$ by “pulling” the subgroup $G_{k}$ via the maps $J_{s,t}$ as far as we can until an element diverges to $\infty$, or we reach time $1$. Note that Lemma 4.8 implies that the map $t\mapsto J_{s,t}(G_{k})\cap\text{PSL}_{2}\mathbb{C}$ is a path in the geometric topology on $\mathcal{D}$ on the interval $[s,t_{k+1}]$. Observe that this path satisfies the hypotheses of Lemma 4.15, so it follows that
$G_{k+1}=J_{t_{k+1},s}(J_{s,t_{k+1}}(G_{k})\cap\text{PSL}_{2}\mathbb{C})$
is a free factor of $G_{k}$. If $J_{s,t_{k+1}}|_{G_{k}}$ is not finite valued, define $H_{k+1}\leq G_{k}$ to be a non-trivial subgroup so that
$G_{k}=G_{k+1}*H_{k+1}$
and note that $J_{s,t}(\psi)\rightarrow\infty$ as $t\rightarrow t_{k+1}$ for all $\psi\in H_{k+1}$. Note that $J_{s,t_{k+1}}|_{G_{k}}$ must not be finite valued if $t_{k+1}\neq 1$.
We now claim that this process must terminate, with $t_{k}=1$ for some $k<\infty$. Note that $\text{rank}(G_{0})=\text{rank}(\Gamma_{s})<\infty$. Additionally, we can see from the construction that $\text{rank}(G_{k+1})<\text{rank}(G_{k})$ for all $k\geq 0$ such that $t_{k+1}<1$, by Grushko’s theorem (see eg. [STA]), which confirms that the process must terminate after finitely many steps.
If $J_{s,1}$ is finite valued, set $n=0$. Else, set $n\geq 1$ to be the largest integer such that $J_{s,t_{n}}|_{G_{n-1}}$ is not finite valued. Then, observe that $J_{s,1}|_{G_{n}}$ is finite valued, and hence it follows from the construction combined with Proposition 4.6 and Lemma 4.5(2) that $G_{n}=J_{1,s}(\Gamma_{1})=H_{s}^{J}$. Since each $G_{k}=G_{k+1}*H_{k+1}$ for $k<n$, we have
$\Gamma_{s}=G_{0}=H_{s}^{J}*H_{1}*...*H_{n}.$
Finally, observe that $J_{s,t_{k}}|_{G_{k}}$ is finite valued for all $k$. Hence, since each $G_{k}$ is finitely generated, Corollary 4.7 combined with Proposition 4.6 and Lemma 4.5 implies that $t_{k}<t_{k+1}$ for all $k$ such that $t_{k}\neq 1$. ∎
4.4 Path Components
In this subsection we will show that $\mathcal{H}_{\infty}$ is not path connected. This contrasts with Theorem A, in which we determined that $\mathcal{H}_{\infty}$ is connected. We will describe an infinite family of hyperbolic $3$-manifolds, which we will call symmetric infinite type $(G,N)$-glued hyperbolic $3$-manifolds, such that for each symmetric infinite type $(G,N)$-glued hyperbolic $3$-manifold $M$, the leaf
$\ell(M)=\{(M,f)\in\mathcal{H}\;|\;f\in\mathcal{F}M\}$
is a path component of $\mathcal{H}_{\infty}$.
Construction of $(G,N)$-glued hyperbolic $3$-manifolds: Let $N$ be a compact, oriented, connected, irreducible, atoroidal, acylindrical $3$-manifold where each of the $02n>$ 0 components of $\partial N$ is incompressible and has negative Euler characteristic. Fix an orientation reversing involution $\tau:\partial N\rightarrow\partial N$ where $\tau$ preserves no component of $\partial N$. Observe that $\tau$ partitions the boundary components of $N$ into $n$ pairs; accordingly, assign each component of $\partial N$ a unique label, either $S_{j}^{-}$ or $S_{j}^{+}$ for $1\leq j\leq n$, so that each $\tau(S_{j}^{-})=S_{j}^{+}$.
Let $G$ be a $2n$-regular directed connected graph with a labeling of the edges by $n$ colors, call the colors $\{1,...,n\}$, so that each vertex is incident to exactly one incoming edge and one outgoing edge of each color. Associate to each vertex of $G$ a copy of $N$, and suppose that $v_{1},v_{2}\in V(G)$ are two vertices of $G$ joined by an edge labeled $j$, directed from $v_{1}$ to $v_{2}$. Then, glue the copy of $S_{j}^{-}$ in the $v_{1}$ copy of $N$ to the copy of $S_{j}^{+}$ in the $v_{2}$ copy of $N$ according to $\tau$. This gluing is depicted in Figure 5 for $j=1$. Let $M$ be the manifold resulting from performing the analogous gluing at each edge of $G$. It follows from a result of Souto and Stover [SOUSTO] that $M$ admits a complete hyperbolic metric. For any hyperbolic metric on $M$, we call $M$ a $(G,N)$-glued hyperbolic $3$-manifold.
Figure 5: This depicts a portion of the construction of a $(G,N)$-glued hyperbolic $3$-manifold, with $G$ overlaying the copies of $N$. The edges of $G$ labeled with a $1$ are depicted as dashed and those labeled with a $2$ are depicted solid.
For $M$ a $(G,N)$-glued hyperbolic $3$-manifold as above, we let $\text{Aut}(G)$ denote the set of graph automorphisms of $G$ which preserve edge colorings and orientations. Each $\varphi\in\text{Aut}(G)$ yields an orientation preserving homeomorphism of $M$ by mapping the copy of $N$ corresponding to each vertex $v\in V(G)$ to the copy of $N$ corresponding to $\varphi(v)$; this map is a well defined homeomorphism, since $\varphi$ preserves the gluing data, so we obtain an injective homomorphism $\text{Aut}(G)\hookrightarrow\text{Homeo}^{+}(M)$ into the group of orientation preserving self homeomorphisms. We say that $M$ is symmetric if $\text{Aut}(G)$ acts transitively on the vertex set $V(G)$, and that $M$ is infinite type if $|V(G)|=\infty$.
Note that the topological conditions on $N$ are quite general. Work of Thurston (see [THUI], [KAP]) and McMullen [MCM] implies that these topological conditions are jointly equivalent to $N$ admitting a hyperbolic metric with totally geodesic boundary, an even number of boundary components, and no rank-$2$ cusps. Additionally, note that Thurston’s work also implies that the compact manifold $N/\tau$ obtained by gluing the boundary components of $N$ as determined by the involution $\tau$ is hyperbolizable, and Mostow rigidity implies that this hyperbolic structure is unique.
There are also many examples of infinite oriented colored graphs $G$ where $\text{Aut}(G)$ acts transitively on $V(G)$, for example an appropriately colored and oriented $2n$-regular tree or $n$-dimensional lattice graph.
In the case where $G$ is a finite graph, $M$ is compact. For such $M$, Mostow rigidity implies that $M$ has a unique hyperbolic structure and we recall from Corollary 3.2 and Theorem A that $\ell(M)$ is a path component of $\mathcal{H}$. The main results of the section are that these statements also hold in the case that $M$ is symmetric and infinite type, in which case $M$ must have infinite volume and be non-tame.
Theorem D.
For any symmetric infinite type $(G,N)$-glued hyperbolic $3$-manifold $M$, the leaf $\ell(M)$ is a path component of $\mathcal{H}_{\infty}$.
We can determine the topology of the path component $\ell(M)$ under the hypotheses of Theorem D. Recall that map $L_{M}:\mathcal{F}M\rightarrow\mathcal{H}$ defined by $L_{M}(f)=(M,f)$ is continuous by Proposition 3.1. As discussed in Section 3.1, $L_{M}$ descends to a continuous bijection $\mathcal{F}M/\text{Isom}^{+}(M)\rightarrow\ell(M)$. Since $M$ is nonelementary, $\text{Isom}^{+}(M)$ acts freely and properly discontinuously on $\mathcal{F}M$, so $\mathcal{F}M/\text{Isom}^{+}(M)$ has a manifold topology inherited from $M$. In the case where $M$ is a symmetric infinite type $(G,N)$-glued hyperbolic, as in Theorem D, Proposition 4.16 (below) implies that $\mathcal{F}M/\text{Isom}^{+}(M)$ is compact. It then follows that the path component $\ell(M)$ with the subspace topology is homeomorphic to the manifold $\mathcal{F}M/\text{Isom}^{+}(M)$.
Theorem D gives the following immediate corollary.
Corollary E.
$\mathcal{H}_{\infty}$ is not path connected.
Before continuing, it is worthwhile contrasting Theorem D with Example B.3, which will highlight the importance of the assumption that the manifold $N$ in the construction of a $(G,N)$-glued hyperbolic $3$-manifold does not have any toroidal boundary components in addition to the boundary components of negative Euler characteristic which are to be glued. Let $S$ be a closed surface of genus at least $2$ and let $C_{0}\subseteq S$ be a disjoint union of simple closed curves on $S$ which fill $S$, giving a pants decomposition. Let $\tau:S\rightarrow S$ be a pseudo-Anosov homeomorphism and set
$C=\bigcup_{n\in\mathbb{Z}}\tau^{n}(C_{0})\times\{n+(1/2)\}\subseteq S\times\mathbb{R}.$
In Example B.3, a path in $\mathcal{H}$ is constructed between a geometrically finite framed hyperbolic $3$-manifold and a framed manifold $M_{\infty}$, which is homeomorphic to $(S\times\mathbb{R})\backslash C$, and has a unique hyperbolic structure.
Let $T_{0}\subseteq S\times[0,1]$ be a small open tubular neighborhood of $C_{0}\times\{1/2\}$ so that the interior of the $N^{\prime}=(S\times[0,1])\backslash T_{0}$ is homeomorphic to $(S\times(0,1))\backslash(C_{0}\times\{1/2\})$. Note that $N^{\prime}$ is a compact, oriented, connected, irreducible, atoroidal, acylindrical $3$-manifold. Let $S_{-}^{\prime}=S\times\{0\}\subseteq N^{\prime}$ and $S_{+}^{\prime}=S\times\{1\}\subseteq N^{\prime}$. For $n\in\mathbb{Z}$, enumerate copies $N_{n}^{\prime}$ of $N^{\prime}$. Now, let $M^{\prime}$ be the result of using the homeomorphism $\tau$ to glue the copy of $S_{-}^{\prime}$ in $N_{n}^{\prime}$ to the copy of $S_{+}^{\prime}$ in $N_{n+1}^{\prime}$ for all $n\in\mathbb{Z}$.
Observe that the interior of $M^{\prime}$ is homeomorphic to the framed hyperbolic $3$-manifold $M_{\infty}$, which Example B.3 and Lemma 3.3 tell us is in the same path component of $\mathcal{H}_{\infty}$ as $(\mathbb{H}^{3},\mathcal{O})$. The construction of $M^{\prime}$ is completely analogous to the construction of a $(G,N)$-glued hyperbolic $3$-manifold, where $G$ is the bi-infinite path graph: the only difference is that $N^{\prime}$ has boundary components with Euler characteristic 00 ($\tau$ also does not induce an involution of $\partial N^{\prime}$ which is non-trivial on the torus boundary components, but the result of gluing with such an involution would have essential tori). Hence, we see that the topological constraint that all boundary components of $N$ have negative Euler characteristic is necessary for Theorem D.
We will need some preliminary results before we can prove Theorem D. The first is a strong rigidity result for symmetric $(G,N)$-glued hyperbolic $3$-manifolds, which also tells us that the topological symmetry of a symmetric $(G,N)$-glued hyperbolic $3$-manifold descends to symmetry of the hyperbolic metric. This is a version of a result to appear in a forthcoming paper of Cremaschi and Yarmola [CREMYAR], and the author thanks them for allowing it to appear here.
Proposition 4.16.
[…]
Proof.
Fix $p\in M$ and consider $\rho\in D(\Gamma)$. Since any hyperbolic manifold is a $K(\pi,1)$ space, there exists a homotopy equivalence $h:\mathbb{H}^{3}/\Gamma\rightarrow\mathbb{H}^{3}/\rho(\Gamma)$ such that $\Psi_{\rho}\circ h_{*}=\rho\circ\Psi$, where $\Psi$ (resp. $\Psi_{\rho}$) is the holonomy representation of $\pi_{1}(M)$ (resp. $\pi_{1}(\mathbb{H}^{3}/\rho(\Gamma))$) determined by $\Gamma$ and a $\Gamma$-lift of $p$ (resp. $\rho(\Gamma)$ and an appropriate $\rho_{n}(\Gamma)$-lift of $h(p)$). A result of Souto and Stover [SOUSTO, Lemma 3.3] implies that $h$ must be homotopic to a homeomorphism, so we may assume that $h$ itself is a homeomorphism. We may also assume without loss of generality that $h$ is orientation preserving: otherwise, replace $\rho$ with $\rho^{\prime}\in D(\Gamma)$ where $\rho^{\prime}$ is a representation so that $\mathbb{H}^{3}/\rho^{\prime}(\Gamma)$ is isometric to $\mathbb{H}^{3}/\rho(\Gamma)$ by an orientation reversing isometry.
Let $N_{0}$ denote one of the copies of $N$ in the construction of $M$ as a $(G,N)$-glued hyperbolic $3$-manifold, and assume $p\in N_{0}$. Let $H=\Psi(\pi_{1}(N_{0},p))\leq\Gamma$ denote the subgroup corresponding to $\pi_{1}(N_{0},p)$ under the holonomy representation. It follows from Canary’s covering theorem [DICKCOVER, Corollary C] that $H$ and $\rho(H)$ must be convex cocompact. Since $h$ is orientation preserving and $h(N_{0})$ is homeomorphic to $N_{0}$ by the preceding paragraph, we see $\rho|_{H}\in GF(H)$. Thus, we have a map $AH(\Gamma)\rightarrow\overline{GF}(H)$ defined by $[\rho]\mapsto[\rho|_{H}]$. By work of Ahlfors, Bers, Marden, and Sullivan (see eg. [MAT]) we may identify $\overline{GF}(H)$ with the Teichmüller space $\mathcal{T}(\partial N_{0})$. By composition, we obtain a map $\chi:AH(\Gamma)\rightarrow\mathcal{T}(\partial N_{0})$. Our goal will be to show that $\chi$ is a constant map, at which point the two claims will follow from the topological symmetry of $M$.
As in the above construction of $(G,N)$-glued hyperbolic $3$-manifolds, enumerate the components of $\partial N_{0}$ by $S_{j}^{-}$ and $S_{j}^{+}$ for $1\leq j\leq n$, so that the involution $\tau$ in the construction satisfies $\tau(S_{j}^{-})=S_{j}^{+}$ for each $j$. Let $N_{j}^{-}$ and $N_{j}^{+}$ denote the copies of $N$ to which $N_{0}$ is glued along $S_{j}^{-}$ and $S_{j}^{+}$, respectively. Note that we may identify
$\mathcal{T}(\partial N_{0})=\prod_{1\leq j\leq n}\mathcal{T}(S_{j}^{-})\times\mathcal{T}(S_{j}^{+}),\;\;\;\;\;\mathcal{T}(\overline{\partial N_{0}})=\prod_{1\leq j\leq n}\mathcal{T}(\overline{S_{j}^{-}})\times\mathcal{T}(\overline{S_{j}^{+}})$
where $\overline{S}$ denotes the surface $S$ with the reversed orientation. From the gluing map $\tau$, we obtain Teichmüller isometries $\tau_{j}^{-}:\mathcal{T}(S_{j}^{-})\rightarrow\mathcal{T}(\overline{S_{j}^{+}})$ and $\tau_{j}^{+}:\mathcal{T}(S_{j}^{+})\rightarrow\mathcal{T}(\overline{S_{j}^{-}})$. Let $Y=\chi(AH(\Gamma))$, and let $D_{j}^{-}$ and $D_{j}^{+}$ denote the projections of $Y$ to the $\mathcal{T}(S_{j}^{-})$ and $\mathcal{T}(S_{j}^{+})$ factors of $\mathcal{T}(\partial N_{0})$, respectively.
Fix $\mu\in D_{j}^{-}$, for some $1\leq j\leq n$. Let $H^{\prime}\leq\Gamma$ be a subgroup representing the conjugacy class corresponding to $\pi_{1}(N_{j}^{+})$ under the holonomy representation. The symmetry of $M$ implies that for some $\rho\in D(\Gamma)$, $\mu$ appears as the conformal boundary of the end of $\mathbb{H}^{3}/\rho(H^{\prime})$ corresponding to the copy of $S_{j}^{-}$ appearing in $N_{j}^{+}$. Let $\sigma_{N_{0}}:\mathcal{T}(\partial N_{0})\rightarrow\mathcal{T}(\overline{\partial N_{0}})$ denote Thurston’s skinning map corresponding to $N_{0}$ (see eg. [KENT]) and observe that $\tau_{j}^{-}(\mu)$ is the $\mathcal{T}(\overline{S_{j}^{+}})$ coordinate of $\sigma_{N_{0}}(\chi(\rho))$, since $N_{0}$ is attached to $N_{j}^{+}$ by gluing the $N_{0}$ copy of $S_{j}^{+}$ to the $N_{j}^{+}$ copy of $S_{j}^{-}$. Therefore, $\tau_{j}^{-}(\mu)$, and hence $\tau_{j}^{-}(D_{j}^{-})$, is contained in the projection of $\sigma_{N_{0}}(Y)$ to $\mathcal{T}(\overline{S_{j}^{+}})$. A similar argument establishes that each $\tau_{j}^{i}(D_{j}^{+})$ is contained in the projection of $\sigma_{N_{0}}(Y)$ to $\mathcal{T}(\overline{S_{j}^{-}})$.
Since $N_{0}$ is acylindrical, the bounded image theorem [KENT] implies that $\sigma_{N_{0}}(Y)$ is bounded. Hence, it follows from the preceding paragraph that each $d_{j}^{-}=\text{diam}_{\mathcal{T}}(D_{j}^{-})$ and $d_{j}^{+}=\text{diam}_{\mathcal{T}}(D_{j}^{+})$ are finite, where $\text{diam}_{\mathcal{T}}$ is the diameter of these subsets in their respective Teichmüller spaces, with respect to the Teichmüller metric. If any $d_{k}^{-}>0$ or $d_{k}^{+}>0$, then McMullen’s strict contracting of the skinning map [MCM] implies that for each $j$,
$d_{j}^{-}<\max\{d_{1}^{-},...,d_{n}^{-},d_{1}^{+},...,d_{n}^{+}\}\;\;\;\;\;\text{ and }\;\;\;\;\;d_{j}^{+}<\max\{d_{1}^{-},...,d_{n}^{-},d_{1}^{+},...,d_{n}^{+}\}$
again by the preceding paragraph. Not all of these inequalities can hold, so we conclude that $d_{j}^{-}=d_{j}^{+}=0$ for all $j$, which completes the proof that $\chi$ is constant. ∎
We will also need a result on discrete extensions of Kleinian groups. Several authors, for example see Susskind [SUSS] and Anderson [ANDER], have examined sets of the form
$\{\psi\in\text{PSL}_{2}\mathbb{C}\;|\;\langle\Gamma,\psi\rangle\text{ is discrete \& }\text{fix}(\psi)\cap\Lambda(\Gamma)\neq\varnothing\},$
where $\text{fix}(\psi)\subseteq S^{2}_{\infty}$ is the set of points fixed by $\psi$, placing various constraints on the Kleinian group $\Gamma$, or additional restrictions on $\psi$. Most of these results qualitatively describe the elements $\psi$ in such a set. Our primary focus will be the cardinality of such a set, though some details on the elements will be gleaned along the way.
For a Kleinian group $\Gamma$, define
$C_{\Gamma}=\{\psi\in\text{PSL}_{2}\mathbb{C}\;|\;\psi\Gamma\psi^{-1}\cap\Gamma\text{ is nonelementary}\}.$
Lemma 4.17.
For any Kleinian group $\Gamma$, $C_{\Gamma}$ is countable.
Proof.
Define a map $F:\Gamma^{4}\rightarrow C_{\Gamma}$ as follows. Fix $\varphi_{1},\varphi_{2},\varphi_{3},\varphi_{4}\in\Gamma$. If $\langle\varphi_{1},\varphi_{2}\rangle$ is nonelementary and there exists $\psi\in\text{PSL}_{2}\mathbb{C}$ such that $\varphi_{3}=\psi\varphi_{1}\psi^{-1}$ and $\varphi_{4}=\psi\varphi_{2}\psi^{-1}$, then Fact 2.6 tells us $\psi$ is uniquely determined, and we define $F(\varphi_{1},\varphi_{2},\varphi_{3},\varphi_{4})=\psi$. Otherwise, define $F(\varphi_{1},\varphi_{2},\varphi_{3},\varphi_{4})=\mathds{1}$. Then, $F$ has countable domain and is surjective, confirming that $C_{\Gamma}$ is countable. ∎
The following proposition will provide the result we need on discrete extensions of Kleinian groups.
Proposition 4.18.
Suppose $\Gamma$ is a purely hyperbolic Kleinian group such that the injectivity radius of $\mathbb{H}^{3}/\Gamma$ is uniformly bounded above, and let
$S=\{\psi\in\text{PSL}_{2}\mathbb{C}\;|\;\langle\Gamma,\psi\rangle\text{ is discrete}\}.$
Then, for each $\psi\in S$, there is some $n\in\mathbb{Z}$ such that $\psi^{n}\in C_{\Gamma}$. In particular, $S$ is countable.
Note that the hypothesis that $\mathbb{H}^{3}/\Gamma$ has injectivity radius bounded above implies that $\Lambda(\Gamma)=S^{2}_{\infty}$, so the fixed point set of every torsion free $\psi\in\text{PSL}_{2}\mathbb{C}$ intersects $\Lambda(\Gamma)$. This result may be of some independent interest due to the lack of any restrictions on $\psi$ in the definition of the set $S$.
Proof of Proposition 4.18.
Fix $\psi\in S$. First, assume that $\text{fix}(\psi)\cap\text{fix}(\varphi)\neq\varnothing$ for some $\varphi\in\Gamma$. Then, the discreteness of $\langle\Gamma,\psi\rangle$ implies that $\langle\psi,\varphi\rangle$ is cyclic, so there is some $n\in\mathbb{Z}$ such that $\psi^{n}\in\langle\varphi\rangle\leq\Gamma\subseteq C_{\Gamma}$.
Now, assume that $\text{fix}(\psi)\cap\text{fix}(\varphi)=\varnothing$ for all $\varphi\in\Gamma$. Note that this implies that for all $\varphi\in\Gamma$ and $n\neq 0$, $\langle\varphi,\psi^{n}\varphi\psi^{-n}\rangle$ is a nonelementary discrete group.
Let $R$ be an upper bound for the injectivity radius of $\mathbb{H}^{3}/\Gamma$ and fix $p\in\mathbb{H}^{3}$. Then, for all $n\in\mathbb{N}$, there exists $\varphi_{n}\in\Gamma$ such that $d_{\mathbb{H}^{3}}(\varphi_{n}\cdot(\psi^{n}\cdot p),\psi^{n}\cdot p)\leq R$, and hence that
$d_{\mathbb{H}^{3}}((\psi^{-n}\varphi_{n}\psi^{n})\cdot p,p)\leq R.$
The set $\{\psi^{-n}\varphi_{n}\psi^{n}\;|\;n\in\mathbb{N}\}$ must be bounded, so the discreteness of $\langle\Gamma,\psi\rangle$ implies that this set is finite. Enumerate
$\{\psi^{-n}\varphi_{n}\psi^{n}\;|\;n\in\mathbb{N}\}=\{a_{1},...,a_{N}\}$
for some $N\geq 1$ and let $S_{i}=\{n\in\mathbb{N}\;|\;\psi^{-n}\varphi_{n}\psi^{n}=a_{i}\}$, for $1\leq i\leq N$.
For each $i$, let $m_{i}=\min S_{i}$. Then, for $n\in S_{i}$, we have
$\begin{split}\psi^{-n}\varphi_{n}\psi^{n}&=a_{i}\\ \varphi_{n}&=\psi^{n}a_{i}\psi^{-n}\\ \varphi_{n}&=\psi^{n-m_{n}}(\psi^{m_{i}}a_{i}\psi^{-m_{i}})\psi^{m_{i}-n}.\end{split}$
Thus, setting $S_{i}^{\prime}=\{n-m_{i}:n\in S_{i}\}$ and renaming $\varphi_{i}^{*}=\psi^{m_{i}}a_{i}\psi^{-m_{i}}=\varphi_{m_{i}}\in\Gamma$, we have $\psi^{n}\varphi_{i}^{*}\psi^{-n}\in\Gamma$ for all $n\in S_{i}^{\prime}$.
Since the sets $S_{1},...,S_{N}$ partition $\mathbb{N}$, van der Waerden’s theorem [VAN] tells us that one of the $S_{i}$, say $S_{1}$, contains an arithmetic progression of length $3$. Then, $S_{1}^{\prime}$ contains an arithmetic progression of length $3$, so there exist $A,D\in\mathbb{N}$ such that $A,A+D,A+2D\in S_{1}^{\prime}$. We may therefore define the subgroups
$\begin{split}H_{1}&=\langle\psi^{A}\varphi_{1}^{*}\psi^{-A},\psi^{A+D}\varphi_{1}^{*}\psi^{-(A+D)}\rangle\leq\Gamma\;\;\;\text{ and}\\ H_{2}&=\langle\psi^{A+D}\varphi_{1}^{*}\psi^{-(A+D)},\psi^{A+2D}\varphi_{1}^{*}\psi^{-(A+2D)}\rangle\leq\Gamma.\end{split}$
As noted, the assumption on $\psi$ implies that each $H_{1},H_{2}$ is nonelementary. We can see that $\psi^{D}H_{1}\psi^{-D}=H_{2}$, so $\psi^{D}\in C_{\Gamma}$, as desired.
Note that for any countable $X\subseteq\text{PSL}_{2}\mathbb{C}$, the set $\{\varphi\in\text{PSL}_{2}\mathbb{C}\;|\;\exists n\in\mathbb{Z}\text{ such that }\varphi^{n}\in X\}$ is countable, so the claim that $S$ is countable now follows from Lemma 4.17. ∎
Finally, we can complete the proof of Theorem D.
Proof of Theorem D.
Let $\Gamma:[0,1]\rightarrow\mathcal{D}$ be a path and write $\Gamma_{t}=\Gamma(t)$ for all $t$. Suppose that $M=\mathbb{H}^{3}/\Gamma_{0}$ is a symmetric infinite type $(G,N)$-glued hyperbolic $3$-manifold. We will show that $\Gamma_{t}$ is conjugate to $\Gamma_{0}$ for all $t$, which will imply the result.
Fix some vertex $v_{0}$ in $G$. Note from the construction of $(G,N)$-glued hyperbolic $3$-manifolds that we can write $M=\bigcup_{n\geq 0}X_{n}$, where each $X_{n}$ is the union of the copies of $N$ which correspond to vertices of $G$ with graph distance at most $n$ from $v_{0}$ (the graph distance being measured disregarding edge orientation). Fix a point $p\in X_{0}$, and for $n\geq 0$ let $H_{n}\leq\Gamma_{0}$ be the subgroup which is the image of $\pi_{1}(X_{n},p)\leq\pi_{1}(M,p)$ under a fixed holonomy representation $\pi_{1}(M,p)\rightarrow\Gamma_{0}$ with respect to $p$ and $\Gamma_{0}$.
For each $n\geq 0$, let
$t_{n}=\sup\{t\in[0,1]\;|\;\text{for all }t^{\prime}\leq t\text{, }J_{0,t^{\prime}}|_{H_{n}}\text{ is finite valued}\}$
where the maps $J_{s,t}:\Gamma_{s}\rightarrow\overline{\text{PSL}_{2}\mathbb{C}}$ are those constructed in Equation 1 with respect to the path $\Gamma$. Since each $H_{n}$ is finitely generated, Corollary 4.7 tells us each $t_{n}>0$. We assume for now that $J_{0,t}|_{H_{0}}$ is in fact finite valued for all $t\in[0,1]$, so $t_{0}=1$.
Suppose, for contradiction, that there exists some $n\geq 1$ such that $t_{n}<1$. Observe that $J_{0,t_{n}}|_{H_{n}}$ must not be finite valued, else Corollary 4.7 combined with Proposition 4.6 would provide a contradiction to the definition of $t_{n}$. For each $t\in[0,t_{n})$, Proposition 4.6 implies that $J_{0,t}|_{H_{n}}$ is an injective homomorphism, so $J_{0,t}|_{H_{n}}\in D(H_{n})$. Let $\{s_{k}\}_{k=1}^{\infty}\subseteq[0,t_{n})$ be a sequence such that $s_{k}\rightarrow t_{n}$ as $k\rightarrow\infty$. Since each $X_{n}$ is acylindrical, Thurston’s “$AH(\text{acylindrical})$ is compact” theorem [THUI] tells us that $\{[J_{0,s_{k}}|_{H_{n}}]\}\subseteq AH(H_{n})$ has a convergent subsequence, which we will denote the same way. Note that $J_{0,s_{k}}|_{H_{0}}\rightarrow J_{0,t_{n}}|_{H_{0}}$ in $D(H_{0})$, as $k\rightarrow\infty$, so since $H_{0}$ is nonelementary, Lemma 2.7 tells us there exists $\rho\in D(H_{n})$ such that $J_{0,s_{k}}|_{H_{n}}\rightarrow\rho$ as $k\rightarrow\infty$. In particular, for all $\varphi\in H_{n}$, $J_{0,s_{k}}(\varphi)$ is bounded as $k\rightarrow\infty$, so it follows that the path $t\mapsto J_{0,t}(\varphi)$ does not diverge to $\infty$ as $t$ increases from 00 to $t_{n}$. This, however, contradicts that $J_{0,t_{n}}|_{H_{n}}$ is not finite valued, by the continuity statement in Lemma 4.5. We therefore conclude that $t_{n}=1$ for all $n$ and that $J_{0,t}$ is injective for all $t\in[0,1]$.
Since each $J_{0,t}$ is injective, $t\mapsto[J_{0,t}]$ is a path in $AH(\Gamma_{0})$. Since $M$ is a symmetric $(G,N)$-glued hyperbolic $3$-manifold, Proposition 4.16 combined with the continuity statement in Fact 2.6 tell us that there exists a path $\psi:[0,1]\rightarrow\text{PSL}_{2}\mathbb{C}$ so that for all $\varphi\in\Gamma_{0}$, $\psi(t)J_{0,t}(\varphi)\psi(t)^{-1}=\varphi$. Now, define a path $\Gamma^{\prime}:[0,1]\rightarrow\mathcal{D}$ by $\Gamma^{\prime}(t)=\psi(t)\Gamma_{t}\psi(t)^{-1}$, and write $\Gamma^{\prime}(t)=\Gamma_{t}^{\prime}$. For $s,t\in[0,1]$, let $J_{s,t}^{\prime}:\Gamma_{s}^{\prime}\rightarrow\overline{\text{PSL}_{2}\mathbb{C}}$ denote the maps defined in Equation 1, now corresponding to the path $\Gamma^{\prime}$; similarly, let $(I_{\varphi}^{s})^{\prime}$ denote the maximal intervals constructed there, corresponding to the path $\Gamma^{\prime}$.
Notice that $J_{0,t}^{\prime}$ is the identity on each element of $\Gamma_{0}^{\prime}=\Gamma_{0}$, and in particular $J_{0,t}^{\prime}(\Gamma_{0}^{\prime})=\Gamma_{0}$ for all $t$. Suppose that there exists $\varphi\in\Gamma_{s}^{\prime}\backslash J_{0,s}^{\prime}(\Gamma_{0}^{\prime})$ for some $s\in[0,1]$. Observe that $0\not\in(I_{\varphi}^{s})^{\prime}$, else $J_{0,s}^{\prime}(J_{s,0}^{\prime}(\varphi))=J_{s,s}(\varphi)=\varphi$ by Lemma 4.5(2), contradicting that $\varphi$ is not in the image of $J_{0,s}^{\prime}$. Thus, Lemma 4.5(1) tells us that the path $J_{s,t}(\varphi)$ diverges to infinity in $\text{PSL}_{2}\mathbb{C}$ as $t$ decreases from $s$ towards 00. In particular, we see that $J_{s,t}(\varphi)$ must take uncountably many values as $t$ decreases. Note that the second statement of Proposition 4.16 implies that the compact copies of $N$ in $\mathbb{H}^{3}/\Gamma_{0}$ are all isometric, so $\mathbb{H}^{3}/\Gamma_{0}$ has bounded injectivity radius; $\Gamma_{0}$ is a purely hyperbolic Kleinian group, so we have obtained a contradiction to Proposition 4.18. Thus, $J_{0,t}^{\prime}(\Gamma_{0}^{\prime})=\Gamma_{0}^{\prime}$ for all $t$, so we have shown that there exists a path $\psi:[0,1]\rightarrow\text{PSL}_{2}\mathbb{C}$ such that $\Gamma_{0}=\psi(t)\Gamma_{t}\psi(t)^{-1}$ for all $t\in[0,1]$.
Finally, we need to address our added assumption that $J_{0,t}|_{H_{0}}$ is finite valued for all $t\in[0,1]$. For a general path $\Gamma:[0,1]\rightarrow\mathcal{D}$ where $M=\mathbb{H}^{3}/\Gamma_{0}$ is a symmetric $(G,N)$-glued hyperbolic $3$-manifold, we must have that $t_{0}>0$, so we have shown that $\Gamma_{t}$ must be conjugate to $\Gamma_{0}$ for $t$ sufficiently close to 00. Let $t_{*}\in[0,1]$ be the supremal value such that for all $t<t_{*}$, $\Gamma_{t}$ is conjugate to $\Gamma_{0}$. Suppose for contradiction that $t_{*}<1$. The symmetry of $M$ described in Proposition 4.16 implies that $\ell(M)\subseteq\mathcal{H}$ is compact, or equivalently that the subspace of $\mathcal{D}$ consisting of Kleinian groups that are conjugate to $\Gamma_{0}$ is compact. Thus, it must be the case that $\Gamma_{t_{*}}$ is conjugate to $\Gamma_{0}$. Hence, $\mathbb{H}^{3}/\Gamma_{t_{*}}$ is a symmetric infinite type $(G,N)$-glued hyperbolic $3$-manifold, so applying the above result to the path with the restricted domain $\Gamma|_{[t_{*},1]}$, we see that $\Gamma_{t}$ must be conjugate to $\Gamma_{t_{*}}$ – and hence to $\Gamma_{0}$ – for $t$ sufficiently close to $t_{*}$. This contradicts the definition of $t_{*}$, so we conclude that $t_{*}=1$, as was to be shown. ∎
[…]